### 18 Mar 2026 v1

# Counting subalgebras of $\mathfrak{o}^{n}$

Abstract.
Let $\mathfrak{o}$ be a compact discrete valuation ring and $n\geq 2$. We introduce a method to study the cotype zeta function of subalgebras of $\mathfrak{o}^{n}$. This multivariable series encodes the number of finite-index subalgebras $\Lambda$ of the $\mathfrak{o}$-algebra $\mathfrak{o}^{n}$ of a given elementary divisor type. We express this zeta function as a finite sum of $\mathfrak{o}$-adic integrals and compute these integrals in many cases.
As a first application, we recover known results in a natural way from our approach. For instance, we obtain a lower bound for the abscissa of convergence of the subalgebra zeta function of $\mathfrak{o}^{n}$ by exhibiting an explicit pole. We also determine the number of irreducible subrings of $\mathfrak{o}^{n}$ of small index.
As a second application, we give an explicit formula for the cotype zeta function of subalgebras of $\mathfrak{o}^{4}$.
Key words and phrases:
cotype zeta function, subring zeta function, $p$-adic integration
1991 Mathematics Subject Classification:
11M41, 11S40, 20E07
1. Introduction
We consider the ring $\mathbb{Z}^{n}$ equipped with component-wise addition and multiplication. For each positive integer $k$, let $f_{n}(k)$ be the number of subrings (= subgroups closed under multiplication) of $\mathbb{Z}^{n}$ of index $k$ containing the identity $(1,\ldots,1)$. The sequence $\big(f_{n}(k)\big)_{k\geq 1}$ and the asymptotic behaviour of its partial sums have attracted considerable attention in recent years because of deep connections with the distribution of orders in number fields; see [9, 1, 7, 8, 13].
A natural generating function for the sequence $\big(f_{n}(k)\big)_{k\geq 1}$ is the zeta function
$\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle 1,<}(s):=\sum_{k=1}^{\infty}\frac{f_{n}(k)}{k^{s}}.$
More generally, let $L$ be a ring additively isomorphic to $\mathbb{Z}^{n}$. It is not necessary to assume that $L$ is commutative nor associative. Following Grunewald, Segal, and Smith [6], one considers the Dirichlet series
$\zeta_{L}^{\scriptscriptstyle<}(s):=\sum_{k=1}^{\infty}\frac{a_{L}^{\scriptscriptstyle<}(k)}{k^{s}},$
where $a_{L}^{\scriptscriptstyle<}(k)$ denotes the number of subrings (= subgroups stable under multiplication) of $L$ of index $k$. When $L$ has an identity one may instead look at the series
$\zeta_{L}^{\scriptscriptstyle 1,<}(s)=\sum_{k=1}^{\infty}\frac{a_{L}^{\scriptscriptstyle 1,<}(k)}{k^{s}}$
enumerating only the finite index subrings containing the identity.
1.1. General properties of subring zeta functions
The Dirichlet series $\zeta_{L}^{\scriptscriptstyle<}(s)$ (or $\zeta_{L}^{\scriptscriptstyle 1,<}(s)$) enjoys the following fundamental properties:
  1. ((a))
    It admits an Euler factorization
    $\zeta_{L}^{\scriptscriptstyle<}(s)=\prod_{p\ \mathrm{prime}}\zeta_{L_{p}}^{\scriptscriptstyle<}(s),$
    where $L_{p}:=L\otimes_{\mathbb{Z}}\mathbb{Z}_{p}$ ($\mathbb{Z}_{p}$ being the ring of $p$-adic integers) and
    $\zeta_{L_{p}}^{\scriptscriptstyle<}(s):=\sum_{\Lambda\leq L_{p}}[L_{p}:\Lambda]^{-s}=\sum_{e=0}^{\infty}\frac{a_{L}^{\scriptscriptstyle<}(p^{e})}{p^{es}},$
    the summation in the middle being over the finite index subrings of $L_{p}$. Each local factor is known to be a rational function in $p^{-s}$ [6].
  2. ((b))
    It has a rational abscissa of convergence, and it extends meromorphically beyond this abscissa [5].
  3. ((c))
    If $\alpha_{L}^{\scriptscriptstyle<}$ is the abscissa of convergence and the continued function has a pole of order $\beta_{L}^{\scriptscriptstyle<}$ at $s=\alpha_{L}^{\scriptscriptstyle<}$, then
    $\sum_{k\leq N}a_{L}^{\scriptscriptstyle<}(k)\sim CN^{\alpha_{L}^{\scriptscriptstyle<}}(\log N)^{\beta_{L}^{\scriptscriptstyle<}-1}\qquad(N\to\infty)$
    for some constant $C>0$ [5].
  4. ((d))
    For almost all primes $p$, the local factor $\zeta_{L_{p}}^{\scriptscriptstyle<}(s)$ satisfies a functional equation under $p\mapsto p^{-1}$; see [16].
1.2. Connection with number fields
When $L=\mathcal{O}_{K}$ is the ring of integers of a number field $K$, the zeta function $\zeta_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(s)$ is also called the order zeta function of $K$, as the orders of $K$ are precisely the finite index subrings of $\mathcal{O}_{K}$ containing the identity.
If $K$ is a quadratic field, one easily finds that $\zeta_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(s)=\zeta(s)$, the Riemann zeta function. When $K$ is a cubic field, Datskovsky and Wright [4] obtained
$\zeta_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(s)=\frac{\zeta_{K}(s)}{\zeta_{K}(2s)}\,\zeta(2s)\,\zeta(3s-1),$
where $\zeta_{K}(s)$ is the Dedekind zeta function of $K$.
When $K$ is a quartic field, Nakanawa [14] computed almost all the local factors of $\zeta_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(s)$. On the other hand, no formula for $\zeta_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(s)$ is known if $K:\mathbb{Q}]\geq 5$. Nevertheless, Kaplan, Marcinek, and Takloo-Bighash [9] showed that for any number field $K$ of degree $\leq 5$, the abscissa of convergence of $\zeta_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(s)$ is 1. They also obtained an expression for the order $\beta_{K}$ of the pole at $s=1$ in terms of the Galois group of the normal closure of $K$. As mentioned above, this suffices to obtain an asymptotic formula of the form
$\sum_{k\leq N}a_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(k)\sim C_{K}N(\log N)^{\beta_{K}-1},$
for some $C_{K}>0$.
If $K:\mathbb{Q}]=n$, then the zeta functions $\zeta_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(s)$ and $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ share the same local factors at the completely ramified primes (since for those primes there is a ring isomorphism $\mathcal{O}_{K}\otimes\mathbb{Z}_{p}=\mathbb{Z}_{p}^{n}$). The relationship was recently strengthened by the following result, originally conjectured in [9]:
Theorem 1.1 ([15, Theorem E]).
If $K$ is a number field of degree $n$, then $\zeta_{\mathcal{O}_{K}}^{\scriptscriptstyle{1,<}}(s)$ and $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ have the same abscissa of convergence.
Thus the analytic behaviour of $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$—in particular its abscissa of convergence—directly informs us about the distribution of orders in degree-$n$ number fields.
1.3. Known explicit formulas
Explicit formulas are known for $n\leq 4$:
$\zeta_{\mathbb{Z}^{2}}^{\scriptscriptstyle{1,<}}(s)=$ $\zeta(s),$
$\zeta_{\mathbb{Z}^{3}}^{\scriptscriptstyle{1,<}}(s)=$ $\frac{\zeta(3s-1)\zeta(s)^{3}}{\zeta(2s)^{2}},$
$\zeta_{\mathbb{Z}^{4}}^{\scriptscriptstyle{1,<}}(s)=$ $\prod_{p}\frac{1}{\bigl(1-p^{-s}\bigr)^{2}\bigl(1-p^{2}p^{-4s}\bigr)\bigl(1-p^{3}p^{-6s}\bigr)}\Bigl(1+4p^{-s}+2p^{-2s}+(4p-3)p^{-3s}+(5p-1)p^{-4s}$
$+(p^{2}-5p)p^{-5s}+(3p^{2}-4p)p^{-6s}-2p^{2}p^{-7s}-4p^{2}p^{-8s}-p^{2}p^{-9s}\Bigr).$
The case $n=2$ is easy; the case $n=3$ is originally due to Datskovsky and Wright [4]; and the case $n=4$ is due to Nakanawa [14]. These cases were later rederived using combinatorial methods by Liu [11]. No explicit formula is known for $n\geq 5$.
In contrast, the subgroup zeta function $\zeta_{\mathbb{Z}^{n}}(s)$ (enumerating all finite index subgroups) is well-understood:
$\zeta_{\mathbb{Z}^{n}}(s)=\zeta(s)\zeta(s-1)\cdots\zeta(s-n+1);$
see [12] for several proofs.
1.4. On the abscissa of convergence
The above formulas for $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ show that $\alpha_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}=1$ for $n\leq 4$. This is also true for $n=5$, while for $n\geq 6$ we have $1\leq\alpha_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}\leq\frac{n}{2}-\frac{7}{6}$; see [9, Theorem 6]. A sharper lower bound for $n\geq 7$ was obtained by Isham [7] following ideas of Brakenhoff [2]:
Theorem 1.2 ([7, Theorem 1.10]).
Let $n\geq 7$. Then
$d_{7}(n):=\max_{0\leq d\leq n-1}\frac{d(n-1-d)}{n-1+d}<\alpha_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}.$
Strictly speaking, Isham’s theorem as originally stated gives the lower bound $d_{7}(n)$ for the abscissa of convergence of each local factor $\zeta_{\mathbb{Z}_{p}^{n}}^{\scriptscriptstyle{1,<}}(s)$. However, the results of [5, Section 4] imply that the abscissa of convergence of the global zeta function $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ is strictly greater than the abscissa of convergence of each of its local factors.
The inequality in the theorem implies that $\alpha_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}>1$ for $n\geq 7$ and $\liminf_{n\to\infty}\frac{\alpha_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}}{n}\geq 3-2\sqrt{2}$.
1.5. Uniformity
We finally mention an important open question concerning the behaviour of the local factors of $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$. Although each $\zeta_{\mathbb{Z}_{p}^{n}}^{\scriptscriptstyle{1,<}}(s)$ is a rational function in $p^{-s}$, it is unknown how this rational expression depends on $p$. We say that $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ is uniform if there exists a rational function $W_{n}(X,Y)\in\mathbb{Q}(X,Y)$ such that
$\zeta_{\mathbb{Z}_{p}^{n}}^{\scriptscriptstyle{1,<}}(s)=W_{n}(p,p^{-s})$
for every prime $p$.
Question 1.3 ([11, Section 7], [17, Question 3.7]).
Is $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ uniform?
The answer is known to be yes for $n\leq 4$ (as the explicit formulas show), but remains open for $n\geq 5$. This question is equivalent to the following:
Question 1.4.
Is $f_{n}(p^{e})$ a polynomial in $p$ for every fixed $e\geq 1$?
The strongest results in this direction are the following.
Theorem 1.5 ([11], [1]).
For every $e\leq 8$, $f_{n}(p^{e})$ is a polynomial in $n$ and $p$.
Mishra and Ray [13] expressed $f_{n}(p^{9})$ as a polynomial in $n$ and $p$ plus a single remaining term $\gamma(n,p)$, which they conjecture is also polynomial.
A key ingredient in these results is the notion of irreducible subrings.
Definition 1.6.
A subring $\Lambda\leq\mathbb{Z}_{p}^{n}$ is called irreducible if every element $(x_{1},\dots,x_{n})\in\Lambda$ satisfies $x_{1}\equiv\cdots\equiv x_{n}\pmod{p}$. We denote by $g_{n}(p^{e})$ the number of irreducible subrings of $\mathbb{Z}_{p}^{n}$ of index $p^{e}$ that contain the identity $(1,\ldots,1)$.
Liu [11] proved that every subring $\Lambda\leq\mathbb{Z}_{p}^{n}$ can be written as a direct product of irreducible subrings $\Lambda_{i}\leq\mathbb{Z}_{p}^{n_{i}}$. Using this decomposition (with notation adjusted to match [1, 7, 8]), he derived the following recursion.
Proposition 1.7 ([11, Proposition 4.4]; see also [8, Proposition 1.7]).
Set $f_{0}(p^{e})=1$ if $e=0$ and $f_{0}(p^{e})=0$ otherwise. Then
$f_{n}(p^{e})=\sum_{i=0}^{e}\sum_{j=1}^{n}\binom{n-1}{j-1}f_{n-j}(p^{e-i})\,g_{j}(p^{i}).$
Hence, computing $f_{n}(p^{e})$ reduces to determining $g_{j}(p^{i})$ for $1\leq j\leq n$ and $i\leq e$. One easily shows that $g_{j}(p^{i})=0$ for i<j1i<j-1 and $g_{j}(p^{j-1})=1$, hence one only needs the values $g_{j}(p^{i})$ for $j\leq\min\{n,e\}$ and $j\leq i\leq e$. This observation also implies that Question 1.4 has a positive answer if the following question does:
Question 1.8.
For $n\leq e$, is $g_{n}(p^{e})$ a polynomial in $p$?
Liu [11, Proposition 4.3] gave a positive answer for $e=n$, Atanasov–Kaplan–Krakoff–Menzel [1, Corollary 3.7] for $e=n+1$, and recently Isham [8, Theorem 1.7] for $e=n+2$.
1.6. From $n$ to n1n-1
By [11, Proposition 2.3], there is an index-preserving correspondence between the lattice of finite index subrings of $\mathbb{Z}_{p}^{n}$ containing the identity $(1,\ldots,1)$ and the lattice of subrings of $\mathbb{Z}_{p}^{n-1}$. This implies that
$\zeta_{\mathbb{Z}_{p}^{n}}^{\scriptscriptstyle{1,<}}(s)=\zeta_{\mathbb{Z}_{p}^{n-1}}^{\scriptscriptstyle{<}}(s)\quad\forall p,\quad\text{and therefore}\quad\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)=\zeta_{\mathbb{Z}^{n-1}}^{\scriptscriptstyle{<}}(s).$
Moreover, in the above correspondence, the irreducible subrings of $\mathbb{Z}_{p}^{n}$ containing the identity correspond to the subrings of $\mathbb{Z}^{n-1}$ included in $p\mathbb{Z}_{p}^{n}$. Hence, we can recover $f_{n}(p^{e})$ and $g_{n}(p^{e})$ respectively as the coefficients of $p^{-es}$ in the power series
$\sum_{\Lambda\leq_{s}\mathbb{Z}_{p}^{n-1}}[\mathbb{Z}_{p}^{n-1}:\Lambda]^{-s}\quad\text{and}\quad\sum_{\begin{subarray}{c}\Lambda\leq_{s}\mathbb{Z}_{p}^{n-1}\\ \Lambda\subseteq p\mathbb{Z}_{p}^{n-1}\end{subarray}}[\mathbb{Z}_{p}^{n-1}:\Lambda]^{-s}.$
1.7. Results
Let $\mathfrak{o}$ be a compact discrete valuation ring with residue field of cardinality $q$. In this paper we develop a new method to study the Dirichlet series
$\Xi_{n}(s)=\sum_{\Lambda\leq_{s}\mathfrak{o}^{n}}[\mathfrak{o}^{n}:\Lambda]^{-s},$
where the sum ranges over finite-index $\mathfrak{o}$-subalgebras of $\mathfrak{o}^{n}$ (i.e. $\mathfrak{o}$-submodules closed under componentwise multiplication). Our ultimate goal is to determine these series for arbitrary $n$. In the present paper we do so for $n=2$ and $n=3$. In the special case $\mathfrak{o}=\mathbb{Z}_{p}$, this recovers the known formulas for
$\zeta_{\mathbb{Z}_{p}^{2}}^{\scriptscriptstyle<}(s)=\zeta_{\mathbb{Z}_{p}^{3}}^{\scriptscriptstyle 1,<}(s)\qquad\text{and}\qquad\zeta_{\mathbb{Z}_{p}^{3}}^{\scriptscriptstyle<}(s)=\zeta_{\mathbb{Z}_{p}^{4}}^{\scriptscriptstyle 1,<}(s).$
Unlike several existing approaches to the computation of $\zeta_{\mathbb{Z}_{p}^{n}}^{\scriptscriptstyle<}(s)$, our method follows more closely the approach introduced by Voll [16] to prove local functional equations for subring zeta functions. Roughly speaking, we introduce multivariable zeta functions
$\Xi_{n,I}\bigl(s_{0},(s_{\iota})_{\iota\in I}\bigr),\qquad I\subseteq[n-1]:=\{1,\ldots,n-1\},$
which keep track of the elementary divisor types of subrings $\Lambda\leq_{s}\mathfrak{o}^{n}$; see §2.3. The series $\Xi_{n}(s)$ is then recovered from these multivariable functions via
$\Xi_{n}(s)=\frac{1}{1-q^{-ns}}+\sum_{\emptyset\neq I\subseteq[n-1]}\Xi_{n,I}\bigl(s,(s)_{\iota\in I}\bigr).$
We compute $\Xi_{n,\{\iota\}}(s_{0},s_{\iota})$ explicitly for arbitrary $n$ and any $1\leq\iota\leq n-1$; see §3. This has several applications.
  1. ((a))
    We obtain the following formulas in a natural way:
    $\#\{\Lambda\leq_{s}\mathfrak{o}^{n}:\ [\mathfrak{o}^{n}:\Lambda]$ $=q^{\,n+1},\ \Lambda\subseteq\pi\mathfrak{o}^{n}\}=\binom{n}{1}_{q},$
    $\#\{\Lambda\leq_{s}\mathfrak{o}^{n}:\ [\mathfrak{o}^{n}:\Lambda]$ $=q^{\,n+2},\ \Lambda\subseteq\pi\mathfrak{o}^{n}\}=\binom{n+1}{2}q^{n-1}+\binom{n}{2}_{q};$
    see §3.2. In the case $\mathfrak{o}=\mathbb{Z}_{p}$, this yields more direct formulas for $g_{n+1}(p^{n+1})$ and $g_{n+1}(p^{n+2})$.
  2. ((b))
    The explicit formula for $\Xi_{n,\{\iota\}}(s,s)$ yields a lower bound for the abscissa of convergence of $\Xi_{n}(s)$. More precisely, for each $1\leq\iota\leq n-1$, we show that the right-most pole of $\Xi_{n,\{\iota\}}(s,s)$ is
    $\max_{1\leq\iota\leq n-1}\frac{\iota(n-\iota)}{n+\iota};$
    in particular, this is a pole of $\Xi_{n}(s)$; see §$3.$ This gives a direct proof of Theorem 1.2.
We also describe a strategy for computing $\Xi_{n,\{i_{1},i_{2}\}}(s_{0},s_{i_{1}},s_{i_{2}})$ for general $n$ and any $\{i_{1},i_{2}\}\subseteq[n-1]$; see §4. We carry out this computation explicitly in the case $n=3$ and $\{i_{1},i_{2}\}=\{1,2\}$; see Proposition 4.2.
The collection of multivariable series $\Xi_{n,I}\bigl(s_{0},(s_{\iota})_{\iota\in I}\bigr)$ encodes the same information as the cotype zeta function of $\mathfrak{o}$-subalgebras of $\mathfrak{o}^{n}$ (or, equivalently, of $\mathfrak{o}$-subalgebras of $\mathfrak{o}^{n+1}$ containing the identity) introduced in [3] and, more generally, in [10]. As a consequence of our computations, we obtain a formula for the cotype zeta function of $\mathfrak{o}$-subalgebras of $\mathfrak{o}^{3}$ (equivalently, of $\mathfrak{o}$-subalgebras of $\mathfrak{o}^{4}$ containing the identity); see Proposition 4.3. This formula was conjectured in [3, Conjecture A2].
Our strategy is to express $\Xi_{n,I}\bigl(s_{0},(s_{\iota})_{\iota\in I}\bigr)$ and its coefficients as sums of integrals over suitable subsets of $\mathfrak{o}\times\mathfrak{o}^{I}\times\operatorname{L}_{I}(\mathfrak{o})$, where $\operatorname{L}_{I}$ is a unipotent group of lower-triangular matrices, multiplied by combinatorial coefficients encoding certain flags over finite fields; see §2.3.
In a forthcoming paper, we apply our method to give a new proof of the formula for $g_{n}(p^{n+2})$ due to Isham [8], and to obtain a formula for $f_{n}(p^{9})$, completing the work of Mishra and Ray [13]. In addition, we present heuristics based on our strategy for computing $\Xi_{n,I}\bigl(s_{0},(s_{\iota})_{\iota\in I}\bigr)$ which lead us to suspect that the abscissa of convergence of $\Xi_{n}(s)$ always lies to the right of $(3-2\sqrt{2})n$.
Notation
$\operatorname{Mat}_{m\times n}(R)$ The set $m\times n$-matrices with coefficients in a ring $R$.
$\operatorname{Mat}_{n}(R)$ $\operatorname{Mat}_{n\times n}(R)$.
$\operatorname{Id}_{n}$ The $n\times n$-identity matrix.
$[n]$ The integers from $1$ to $n$.
$\{i_{1},\ldots,i_{l}\}_{<}$ A set of positive integers with $i_{1}<i_{2}<\cdots<i_{l}$.
$\binom{n}{r}_{X}$ The Gaussian polynomial $\prod_{i=0}^{r-1}\frac{(1-X^{n-i})}{(1-X^{r-i})}$.
#S\#S The cardinality of a set $S$.
2. Local zeta functions as integrals
Throughout this section, $\mathfrak{o}$ denotes a compact discrete valuation ring, $\mathfrak{m}$ its maximal ideal, and $\pi$ a uniformizing parameter, i.e. an element such that $\mathfrak{m}=\pi\mathfrak{o}$. We write $q=\#(\mathfrak{o}/\mathfrak{m})$. For $x$ in the fraction field of $\mathfrak{o}$, we define $|x|=q^{-v(x)}$, where $v$ denotes the $\mathfrak{m}$-adic valuation.
Let $n\in\mathbb{N}$. We consider the $\mathfrak{o}$-module $\mathfrak{o}^{n}$, endowed with componentwise multiplication. For a submodule $\Lambda\leq\mathfrak{o}^{n}$, we write $\Lambda\leq_{s}\mathfrak{o}^{n}$ to indicate that $\Lambda$ is a finite-index submodule which is closed under multiplication.
In this section, following the approach of [16], we derive an expression for the generating series
$\sum_{\Lambda\leq_{s}\mathfrak{o}^{n}}[\mathfrak{o}^{n}:\Lambda]^{-s}$
as a sum of integrals reflecting the elementary-divisor types of the submodules.
In §2.1, we parametrize submodules of $\mathfrak{o}^{n}$ by lower-triangular matrices according to their elementary-divisor types. In §2.2, we review some basic facts about integration, and in §2.3, we present the integral expression for the above series.
In this section, we use the following special notation. For a matrix $A\in\operatorname{Mat}_{m\times n}(R)$, where $R$ is a commutative ring, we denote by $\langle A\rangle$ the $R$-submodule generated by the row vectors of $A$.
2.1. Parameterization of finite-index submodules
Any submodule $\Lambda\leq\mathfrak{o}^{n}$ is of the form $\Lambda=\langle A\rangle$ for some $A\in\operatorname{Mat}_{n}(\mathfrak{o})$. The submodule $\Lambda$ has finite index if and only if $A$ is nonsingular, i.e. $\det(A)\neq 0$. If $A,B\in\operatorname{Mat}_{n}(\mathfrak{o})$ are nonsingular, then $\langle A\rangle=\langle B\rangle$ if and only if $A=UB$ for some $U\in\operatorname{GL}_{n}(\mathfrak{o})$.
Let $\Lambda\leq\mathfrak{o}^{n}$ be a finite-index submodule. By the elementary divisor theorem, either $\Lambda=\pi^{r_{0}}\mathfrak{o}^{n}$ for some $r_{0}\in\mathbb{N}_{0}$, or there exist a non-empty subset $I=\{i_{1},\ldots,i_{\ell}\}_{<}\subseteq[n-1]$ and $(r_{0},\mathbf{r})\in\mathbb{N}_{0}\times\mathbb{N}^{I}$ such that
$\Lambda=\bigl\langle\pi^{r_{0}}D(I,\mathbf{r})A\bigr\rangle\qquad\text{for some }A\in\operatorname{GL}_{n}(\mathfrak{o}),$
where
$D(I,\mathbf{r})=\operatorname{diag}\!\bigl(\underbrace{\underbrace{\underbrace{\pi^{\sum_{j=1}^{\ell}r_{i_{j}}},\ldots,\pi^{\sum_{j=1}^{\ell}r_{i_{j}}}}_{i_{1}},\pi^{\sum_{j=2}^{\ell}r_{i_{j}}},\ldots,\pi^{\sum_{j=2}^{\ell}r_{i_{j}}}}_{i_{2}},\ldots,\pi^{r_{i_{\ell}}},\ldots,\pi^{r_{i_{\ell}}}}_{i_{\ell}},1,\ldots,1\bigr).$
The index of $\Lambda$ is given by
(1) $\mathfrak{o}^{n}:\Lambda]=q^{nr_{0}+\sum_{\iota\in I}r_{\iota}\,\iota}.$
We write $\nu_{0}(\Lambda):=(I,r_{0},\mathbf{r})$.
Fix $\emptyset\neq I\subseteq[n-1]$ and $(\mathbf{r},r_{0})\in\mathbb{N}^{I}\times\mathbb{N}_{0}$. Our goal is to parametrize the set
$\{\Lambda\leq\mathfrak{o}^{n}:\ \nu_{0}(\Lambda)=(I,r_{0},\mathbf{r})\}$
by suitable lower-triangular matrices.
Consider the surjective map
$\operatorname{GL}_{n}(\mathfrak{o})\to\{\Lambda\leq\mathfrak{o}^{n}:\ \nu_{0}(\Lambda)=(I,r_{0},\mathbf{r})\},\qquad A\mapsto\bigl\langle\pi^{r_{0}}D(I,\mathbf{r})A\bigr\rangle.$
Given $A,B\in\operatorname{GL}_{n}(\mathfrak{o})$, we have
$\langle\pi^{r_{0}}D(I,\mathbf{r})A\rangle=\langle\pi^{r_{0}}D(I,\mathbf{r})B\rangle$
if and only if $A=UB$ for some
$U\in\operatorname{GL}_{n}(\mathfrak{o})\cap D(I,\mathbf{r})^{-1}\operatorname{GL}_{n}(\mathfrak{o})D(I,\mathbf{r}).$
It is straightforward to verify that this intersection is the subgroup
$\Gamma_{I,\mathbf{r}}:=\left\{\begin{pmatrix}U_{i_{1}}&*&\cdots&*&*\\ \pi^{r_{i_{1}}}*&U_{i_{2}-i_{1}}&\ddots&\vdots&*\\ \pi^{r_{i_{1}}+r_{i_{2}}}*&\pi^{r_{i_{2}}}*&\ddots&*&*\\ \vdots&\vdots&\ddots&U_{i_{\ell}-i_{\ell-1}}&*\\ \pi^{r_{i_{1}}+\cdots+r_{i_{\ell}}}*&\pi^{r_{i_{2}}+\cdots+r_{i_{\ell}}}*&\cdots&\pi^{r_{i_{\ell}}}*&U_{n-i_{\ell}}\end{pmatrix}\right\}$,
where $U_{s}$ denotes an element of $\operatorname{GL}_{s}(\mathfrak{o})$ and * denotes a block matrix of appropriate size with entries in $\mathfrak{o}$. Thus:
Proposition 2.1.
The map $\Gamma_{I,\mathbf{r}}A\mapsto\langle\pi^{r_{0}}D(I,\mathbf{r})A\rangle$ induces a bijection
[…]
$\operatorname{B}_{I}(\mathfrak{o})=\left\{\begin{pmatrix}U_{i_{1}}&*&\cdots&*&*\\ 0&U_{i_{2}-i_{1}}&\ddots&\vdots&*\\ 0&0&\ddots&*&*\\ \vdots&\vdots&\ddots&U_{i_{\ell}-i_{\ell-1}}&*\\ 0&0&\cdots&0&U_{n-i_{\ell}}\end{pmatrix}\right\}$,
and the second is
$\operatorname{L}_{I,\mathbf{r}}:=\left\{\begin{pmatrix}\operatorname{Id}_{i_{1}}&0&\cdots&0&0\\ \pi^{r_{i_{1}}}*&\operatorname{Id}_{i_{2}-i_{1}}&\ddots&\vdots&0\\ \pi^{r_{i_{1}}+r_{i_{2}}}*&\pi^{r_{i_{2}}}*&\ddots&0&0\\ \vdots&\vdots&\ddots&\operatorname{Id}_{i_{\ell}-i_{\ell-1}}&0\\ \pi^{r_{i_{1}}+\cdots+r_{i_{\ell}}}*&\pi^{r_{i_{2}}+\cdots+r_{i_{\ell}}}*&\cdots&\pi^{r_{i_{\ell}}}*&\operatorname{Id}_{n-i_{\ell}}\end{pmatrix}\right\}$.
[…]
$R^{n}=\Lambda_{0}\supset\Lambda_{i_{1}}\supset\cdots\supset\Lambda_{i_{\ell}}\supset\Lambda_{n}=0$
such that $R^{n}/\Lambda_{\iota}$ is free of rank $\iota$ for all $\iota\in I$. In particular, each successive quotient $\Lambda_{i_{j-1}}/\Lambda_{i_{j}}$ is free.
The natural action of $\operatorname{GL}_{n}(R)$ on $R^{n}$ by right multiplication induces a transitive action on the set of $I$-flags. Let $e_{1},\ldots,e_{n}$ be the standard basis of $R^{n}$, and consider the canonical $I$-flag
$R^{n}=\Delta_{0}\supset\Delta_{i_{1}}\supset\cdots\supset\Delta_{i_{\ell}}\supset\Delta_{n}=0,$
where $\Delta_{\iota}=\langle e_{j}:j>\iota\rangle$ for $\iota\in I$. If $A\in\operatorname{GL}_{n}(R)$ has row vectors $A_{1},\ldots,A_{n}$, then the image of the canonical flag under $A$ is the flag
$\langle A\rangle_{I}\,:\,R^{n}\supset\langle A_{i_{1}+1},\ldots,A_{n}\rangle\supset\cdots\supset\langle A_{i_{\ell}+1},\ldots,A_{n}\rangle\supset 0.$
[…]
Definition 2.2.
A matrix $A\in\operatorname{GL}_{n}(R)$ is called $I$-reduced if it satisfies the following conditions:
  1. ((a))
    In each row, the leftmost invertible entry is $1$. We call these entries the leading $1$’s.
  2. ((b))
    All entries above a leading $1$ are zero.
  3. ((c))
    Setting $i_{0}=0$ and $i_{\ell+1}=n$, for each $j=0,\ldots,\ell$, the leading $1$’s in rows $i_{j}+1,\ldots,i_{j+1}$ occur in strictly increasing column positions.
We denote the set of $I$-reduced matrices by $\mathcal{R}_{I}(R)$.
Example 2.3.
The following are examples of $I$-reduced matrices for $n=6$ and $I=\{2,4\}$. The leading $1$’s are displayed in bold.
(a) $R=\mathbb{F}_{p},\quad A=\begin{pmatrix}0&0&0&\mathbf{1}&0&0\\ 0&0&0&0&\mathbf{1}&0\\ \mathbf{1}&0&0&1&-1&0\\ 0&0&0&0&0&\mathbf{1}\\ 0&\mathbf{1}&0&3&1&1\\ 0&0&\mathbf{1}&1&1&0\end{pmatrix};\qquad\text{(b)}\quad R=\mathbb{Z}_{p},\quad A=\begin{pmatrix}0&0&0&\mathbf{1}&0&0\\ 0&0&0&0&\mathbf{1}&0\\ \mathbf{1}&0&0&1&p&0\\ 0&0&0&-p&p&\mathbf{1}\\ p&\mathbf{1}&0&3&1&1\\ p^{2}&0&\mathbf{1}&1&1&p\end{pmatrix}$.
Remark 2.4.
Reduction modulo the maximal ideal of $R$ sends an $I$-reduced matrix over $R$ to an $I$-reduced matrix over the residue field.
Proposition 2.5.
The map
$A\longmapsto\operatorname{B}_{I}(R)A$
[…]
Proof.
Equivalently, every $I$-flag in $R^{n}$ is of the form $\langle A\rangle_{I}$ for a unique $A\in\mathcal{R}_{I}(R)$. We prove this by induction on $n$.
If $n=1$, then $I=\emptyset$ and the claim is immediate. Assume $n>1$. If $I=\emptyset$, the claim is again immediate, so we assume $I\neq\emptyset$ and write $I=\{i_{1}<\cdots<i_{\ell}\}\subseteq[n-1]$.
Let
$R^{n}=\Lambda_{0}\supset\Lambda_{i_{1}}\supset\cdots\supset\Lambda_{i_{\ell}}\supset\Lambda_{n}=0$
be an $I$-flag. We construct a matrix $A\in\mathcal{R}_{I}(R)$ such that $\langle A\rangle_{I}$ equals this flag.
Choose a basis $\{v_{i_{\ell}+1},\ldots,v_{n}\}\subset R^{n}$ for $\Lambda_{i_{\ell}}$ and apply Gaussian elimination to the matrix with row vectors $v_{i_{\ell}+1},\ldots,v_{n}$. Over a local ring, the only difference from the field case is that one must pivot on invertible entries rather than merely nonzero ones. The resulting matrix has the property that the leftmost invertible entry in each row is $1$, the column positions of these leading $1$’s increase from top to bottom, and all entries above each leading $1$ vanish.
Let $A_{i_{\ell}+1},\ldots,A_{n}$ be the row vectors of this matrix. Extend them to a basis
$\{v_{1},\ldots,v_{i_{\ell}},A_{i_{\ell}+1},\ldots,A_{n}\}$
of $R^{n}$ such that, for each $\iota\in I\setminus\{i_{\ell}\}$, the subset $\{v_{\iota+1},\ldots,v_{i_{\ell}},A_{i_{\ell}+1},\ldots,A_{n}\}$ is a basis of $\Lambda_{\iota}$. We may further assume that each $v_{j}$ has zeros in the coordinates corresponding to the leading $1$’s of $A_{i_{\ell}+1},\ldots,A_{n}$. Deleting these coordinates yields vectors $v_{1}^{\prime},\ldots,v_{i_{\ell}}^{\prime}\in R^{i_{\ell}}$.
Let $I^{\prime}:=I\setminus\{i_{\ell}\}$. By the induction hypothesis, there exists an $I^{\prime}$-reduced matrix $A^{\prime}\in\operatorname{GL}_{i_{\ell}}(R)$ whose associated $I^{\prime}$-flag coincides with the flag generated by $v_{1}^{\prime},\ldots,v_{i_{\ell}}^{\prime}$. Let $A_{1}^{\prime},\ldots,A_{i_{\ell}}^{\prime}$ be the rows of $A^{\prime}$, and lift them to vectors $A_{1},\ldots,A_{i_{\ell}}\in R^{n}$ by reinserting zeros in the deleted coordinates. Then the matrix $A$ with rows $A_{1},\ldots,A_{n}$ is $I$-reduced and satisfies $\langle A\rangle_{I}=\{\Lambda_{\iota}\}$.
We now prove uniqueness. Suppose $A,B\in\mathcal{R}_{I}(R)$ satisfy $\langle A\rangle_{I}=\langle B\rangle_{I}$. Let $A^{\prime\prime}$ (resp. $B^{\prime\prime}$) denote the submatrix consisting of the last $n-i_{\ell}$ rows of $A$ (resp. $B$). Then $\langle A^{\prime\prime}\rangle=\langle B^{\prime\prime}\rangle$, so $A^{\prime\prime}=UB^{\prime\prime}$ for some $U\in\operatorname{GL}_{n-i_{\ell}}(R)$.
Reducing modulo the maximal ideal of $R$, the matrices $\overline{A^{\prime\prime}}$ and $\overline{B^{\prime\prime}}$ are row-equivalent matrices in reduced row echelon form over the residue field, and hence coincide. In particular, the leading $1$’s in $A^{\prime\prime}$ and $B^{\prime\prime}$ occur in the same column positions. Restricting the equality $A^{\prime\prime}=UB^{\prime\prime}$ to these pivot columns yields $U=\operatorname{Id}$, and therefore $A^{\prime\prime}=B^{\prime\prime}$.
Finally, let $A^{\prime}$ (resp. $B^{\prime}$) be obtained from $A$ (resp. $B$) by deleting the last $n-i_{\ell}$ rows and the columns containing the leading $1$’s of those rows. Then $A^{\prime}$ and $B^{\prime}$ are $I^{\prime}$-reduced matrices in $\operatorname{GL}_{i_{\ell}}(R)$ and satisfy $\langle A^{\prime}\rangle_{I^{\prime}}=\langle B^{\prime}\rangle_{I^{\prime}}$. By induction, $A^{\prime}=B^{\prime}$, and hence $A=B$. ∎
Remark 2.6.
The proof yields a bijection between $\mathcal{R}_{I}(R)$ and the set of $I$-flags in $R^{n}$. In particular, for $R=\mathbb{F}_{q}$ we obtain
$\mathcal{R}_{I}(\mathbb{F}_{q})=\binom{n}{I}_{q}.$
We now return to our original problem. For $I\subseteq[n-1]$ and $\mathbf{r}\in\mathbb{N}^{I}$, recall that we defined a subgroup $\Gamma_{I,\mathbf{r}}\subseteq\operatorname{GL}_{n}(\mathfrak{o})$ admitting a decomposition
$\Gamma_{I,\mathbf{r}}=\operatorname{B}_{I}(\mathfrak{o})\cdot\operatorname{L}_{I,\mathbf{r}}.$
[…]
Proposition 2.7.
The map
[…]
is a bijection.
Proof.
By Propositions 2.5 and 2.1, we obtain a surjective map
[…]
If $A,B\in\mathcal{R}_{I}(\mathfrak{o})$ have the same image, then $A=UB$ for some $U\in\Gamma_{I,\mathbf{r}}$. Writing $U=U^{\prime}U^{\prime\prime}$ with $U^{\prime}\in\operatorname{B}_{I}(\mathfrak{o})$ and $U^{\prime\prime}\in\operatorname{L}_{I,\mathbf{r}}$, we see that $U^{\prime\prime}B\in\mathcal{R}_{I}(\mathfrak{o})$ and $A=U^{\prime}(U^{\prime\prime}B)\in\mathcal{R}_{I}(\mathfrak{o})$. By Proposition 2.5, this forces $U^{\prime}=\operatorname{Id}$, and hence $A=U^{\prime\prime}B$. Thus $\operatorname{L}_{I,\mathbf{r}}A=\operatorname{L}_{I,\mathbf{r}}B$, as claimed. ∎
Finally, consider the algebraic group $\operatorname{L}_{I}$ consisting of matrices of the form
$\begin{pmatrix}\operatorname{Id}_{i_{1}}&0&\cdots&0&0\\ *&\operatorname{Id}_{i_{2}-i_{1}}&\ddots&\vdots&0\\ \vdots&\ddots&\ddots&0&0\\ *&\cdots&*&\operatorname{Id}_{i_{\ell}-i_{\ell-1}}&0\\ *&\cdots&*&*&\operatorname{Id}_{n-i_{\ell}}\end{pmatrix}$.
Let $A\in\mathcal{R}_{I}(\mathfrak{o})$ (or $A\in\mathcal{R}_{I}(\mathfrak{o}/\mathfrak{m})$). Then $A$ can be written uniquely as $A=B\sigma$ with $B\in\operatorname{L}_{I}(\mathfrak{o})$ and $\sigma\in\mathbb{S}_{n}$, where $B\sigma$ denotes the matrix obtained from $B$ by permuting its columns according to $\sigma$. Conversely, given $B\in\operatorname{L}_{I}(\mathfrak{o})$ and $\sigma\in\mathbb{S}_{n}$, the matrix $B\sigma$ is $I$-reduced if and only if $\overline{B}\sigma$ is $I$-reduced, where $\overline{(\cdot)}$ denotes reduction modulo $\mathfrak{m}$.
Definition 2.8.
For $E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})$ we define
$\mathcal{T}_{I,E}:=\{\sigma\in\mathbb{S}_{n}:\ E\sigma\in\mathcal{R}_{I}(\mathfrak{o}/\mathfrak{m})\}.$
Combining these observations with Proposition 2.7, we obtain the following result.
Proposition 2.9.
For each $E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})$ and each $\sigma\in\mathcal{T}_{I,E}$, the map
[…]
is injective. Moreover, as $E$ varies in $\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})$ and $\sigma$ varies in $\mathcal{T}_{I,E}$, the images of these maps are disjoint and together form a partition of $\{\Lambda\leq\mathfrak{o}^{n}:\ \nu_{0}(\Lambda)=(I,r_{0},\mathbf{r})\}$.
2.2. Preliminaries on $\pi$-adic integration
Every compact topological group admits a (left) Haar measure. In this subsection, we briefly recall the basic facts about $\pi$-adic integration that will be used throughout the paper. We will mainly be concerned with two classes of examples.
The first example is the additive group $\mathfrak{o}^{N}$ for some $N\in\mathbb{N}$. In this case, the normalized Haar measure is the unique Borel measure $\mu$ on $\mathfrak{o}^{N}$ satisfying $\mu(\mathfrak{o}^{N})=1$ and $\mu(x+\mathcal{A})=\mu(\mathcal{A})$ for every Borel set $\mathcal{A}\subseteq\mathfrak{o}^{N}$ and every $x\in\mathfrak{o}^{N}$. For $N=1$, this definition implies that
$\mu(a\mathfrak{o})=|a|\quad\text{and}\quad\mu\bigl(\{x\in\mathfrak{o}:\ |x|=q^{-k}\}\bigr)=(1-q^{-1})q^{-k}\qquad(k\geq 0).$
The normalized Haar measure on $\mathfrak{o}^{N}$ is the product measure associated with the normalized Haar measure on $\mathfrak{o}$. When integrating a complex-valued function $f(x_{1},\ldots,x_{N})$ over an open subset $U\subseteq\mathfrak{o}^{N}$ with respect to this measure, we write
$\int_{U}f(x_{1},\ldots,x_{N})\,|dx_{1}\cdots dx_{N}|\qquad\text{or simply}\qquad\int_{U}f(\mathbf{x})\,|d\mathbf{x}|.$
We will use Haar measure to reinterpret power series in $q^{-s}$ (or in several variables) as integrals.
Example 2.10.
The geometric series $\sum_{k=0}^{\infty}q^{-ks}$ can be expressed as an integral over $\mathfrak{o}$ as follows. For $k\geq 0$ we have
$q^{-ks}$ $=q^{-ks}\,\frac{1}{(1-q^{-1})q^{-k}}\int_{|x|=q^{-k}}|dx|$
$=\frac{1}{1-q^{-1}}\int_{|x|=q^{-k}}|x|^{s-1}\,|dx|.$
Summing over all $k\geq 0$ yields
$\sum_{k=0}^{\infty}q^{-ks}=\frac{1}{1-q^{-1}}\int_{\mathfrak{o}}|x|^{s-1}\,|dx|.$
More generally, for power series of the form $\sum_{k=0}^{\infty}a_{k}q^{-ks}$, the coefficients $a_{k}$ can often be interpreted as measures of suitable subsets of $\mathfrak{o}$ or of $\mathfrak{o}^{N}$ for some $N>1$. Integral representations of this kind are particularly useful, as they allow one to partition domains of integration and apply the change-of-variables formula. Specifically, if $f_{1},\ldots,f_{N}$ are analytic functions defining a bijection between open subsets $\mathcal{A},\mathcal{B}\subseteq\mathfrak{o}^{N}$, then for any continuous function $h:\mathcal{B}\to\mathbb{C}$ one has
$\int_{\mathcal{B}}h(y_{1},\ldots,y_{N})\,|dy_{1}\cdots dy_{N}|=\int_{\mathcal{A}}h\bigl(f_{1}(\mathbf{x}),\ldots,f_{N}(\mathbf{x})\bigr)\,\bigl|\det J(f_{1},\ldots,f_{N})(\mathbf{x})\bigr|\,|dx_{1}\cdots dx_{N}|.$
Our second example of a compact topological group is the multiplicative group $\operatorname{L}_{I}(\mathfrak{o})$ associated with a subset $I\subseteq[n-1]$. Its normalized Haar measure is the unique Borel measure $\mu$ satisfying $\mu\big(\operatorname{L}_{I}(\mathfrak{o})\big)=1$ and $\mu(\mathcal{A}B)=\mu(\mathcal{A})$ for every Borel set $\mathcal{A}\subseteq\operatorname{L}_{I}(\mathfrak{o})$ and every $B\in\operatorname{L}_{I}(\mathfrak{o})$. In particular, if $\Gamma\subseteq\operatorname{L}_{I}(\mathfrak{o})$ is an open subgroup and $\mathcal{A}\subseteq\operatorname{L}_{I}(\mathfrak{o})$ satisfies $\Gamma\mathcal{A}=\mathcal{A}$, then
(2) $(\Gamma\backslash\mathcal{A})=\frac{\mu(\mathcal{A})}{\mu(\Gamma)}.$
For our applications, we will need to compute the measure of the subgroup $\Gamma=\operatorname{L}_{I,\mathbf{r}}$, defined in §2.1. This computation is simplified by the following observation. The group $\operatorname{L}_{I}(\mathfrak{o})$ can be identified, as a topological space, with $\mathfrak{o}^{N}$, where
(3) $N=i_{1}(i_{2}-i_{1})+i_{2}(i_{3}-i_{2})+\cdots+i_{\ell-1}(i_{\ell}-i_{\ell-1})+i_{\ell}(n-i_{\ell}).$
Let $\mu^{\prime}$ denote the measure on $\operatorname{L}_{I}(\mathfrak{o})$ induced by the Haar measure on $\mathfrak{o}^{N}$ under this identification. We claim that $\mu^{\prime}=\mu$.
Indeed, $\mu^{\prime}\big(\operatorname{L}_{I}(\mathfrak{o})\big)=1$ since $\mathfrak{o}^{N}$ has measure $1$. Fix $B\in\operatorname{L}_{I}(\mathfrak{o})$. Right multiplication by $B$ corresponds to an affine transformation $\varphi:\mathfrak{o}^{N}\to\mathfrak{o}^{N}$ of the form $\varphi(\mathbf{x})=\lambda+T\mathbf{x}$, with $\lambda\in\mathfrak{o}^{N}$ and $T\in\operatorname{Mat}_{N}(\mathfrak{o})$. Since $\varphi$ is bijective, we have $T\in\operatorname{GL}_{N}(\mathfrak{o})$, and hence $\det(T)$ is a unit in $\mathfrak{o}$. Therefore $|\det(T)|=1$, and it follows that
$\mu^{\prime}(\mathcal{A}B)=\mu^{\prime}\big(\varphi(\mathcal{A})\big)=|\det(T)|\,\mu^{\prime}(\mathcal{A})=\mu^{\prime}(\mathcal{A}).$
By uniqueness of Haar measure, this implies $\mu^{\prime}=\mu$.
Using this identification, one readily obtains
(4) $\mu\bigl(\operatorname{L}_{I,\mathbf{r}}\bigr)=q^{-\sum_{\iota\in I}r_{\iota}\,\iota(n-\iota)}.$
2.3. Representations in terms of $\pi$-adic integrals
We continue to work with the notation and assumptions introduced in the previous sections. Recall that we write $\Lambda\leq_{s}\mathfrak{o}^{n}$ to indicate that $\Lambda$ is a finite-index submodule of $\mathfrak{o}^{n}$ which is closed under componentwise multiplication.
Our main objective is to compute the generating series
(5) $\Xi_{n}(s):=\sum_{\Lambda\leq_{s}\mathfrak{o}^{n}}[\mathfrak{o}^{n}:\Lambda]^{-s}.$
For $n=1$, this series is elementary:
$\Xi_{1}(s)=\sum_{r_{0}=0}^{\infty}q^{-r_{0}s}=\frac{1}{1-q^{-s}}.$
Henceforth we assume that $n\geq 2$.
Using the parametrization of finite-index submodules described in §2.1, we may write
$\Xi_{n}(s)$ $=\sum_{r_{0}=0}^{\infty}[\mathfrak{o}^{n}:\pi^{r_{0}}\mathfrak{o}^{n}]^{-s}+\sum_{\emptyset\neq I\subseteq[n-1]}\sum_{(r_{0},\mathbf{r})\in\mathbb{N}_{0}\times\mathbb{N}^{I}}q^{-snr_{0}-\sum_{\iota\in I}sr_{\iota}\,\iota}\,h_{n}(I,r_{0},\mathbf{r}),$
where
$h_{n}(I,r_{0},\mathbf{r}):=\#\bigl\{\Lambda\leq_{s}\mathfrak{o}^{n}:\ \nu_{0}(\Lambda)=(I,r_{0},\mathbf{r})\bigr\}.$
In order to keep track of the elementary-divisor types of the subrings, it is convenient to refine this construction. For each non-empty subset $I\subseteq[n-1]$, we therefore introduce the multivariable generating series
$\Xi_{n,I}\bigl(s_{0},(s_{\iota})_{\iota\in I}\bigr)$ $=\sum_{(r_{0},\mathbf{r})\in\mathbb{N}_{0}\times\mathbb{N}^{I}}q^{-s_{0}nr_{0}-\sum_{\iota\in I}r_{\iota}\,\iota\,s_{\iota}}\,h_{n}(I,r_{0},\mathbf{r}).$
With this notation, we obtain the decomposition
(6) $\Xi_{n}(s)=\frac{1}{1-q^{-ns}}+\sum_{\emptyset\neq I\subseteq[n-1]}\Xi_{n,I}\bigl(s,(s)_{\iota\in I}\bigr).$
We first express the quantities $h_{n}(I,r_{0},\mathbf{r})$ in terms of $\pi$-adic integrals.
Proposition 2.11.
For all non-empty $I\subseteq[n-1]$ and $(r_{0},\mathbf{r})\in\mathbb{N}_{0}\times\mathbb{N}^{I}$, we have
$h_{n}(I,r_{0},\mathbf{r})=q^{\sum_{\iota\in I}r_{\iota}\,\iota(n-\iota)}\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}\int\limits_{\begin{subarray}{c}B\in\mathscr{B}_{I,r_{0},\mathbf{r}}\\ \overline{B}=E\end{subarray}}|dB|,$
where
(7) $\tau_{I,E}:=\#\mathcal{T}_{I,E}=\#\{\sigma\in\mathbb{S}_{n}:\ E\sigma\in\mathcal{R}_{I}(\mathfrak{o}/\mathfrak{m})\},$
and $\mathscr{B}_{I,r_{0},\mathbf{r}}\subseteq\operatorname{L}_{I}(\mathfrak{o})$ is the subset defined by the divisibility conditions
$\prod_{\begin{subarray}{c}\iota\in I\\ j\leq\iota<i\end{subarray}}\pi^{r_{\iota}}\;\Big|\;\pi^{r_{0}}\left(\prod_{\begin{subarray}{c}\iota\in I\\ i^{\prime}\leq\iota\end{subarray}}\pi^{r_{\iota}}\right)\left(\sum_{k=1}^{i}B_{ik}B_{i^{\prime}k}(B^{-1})_{kj}\right),\qquad 1\leq j<i\leq i^{\prime}\leq n.$
Proof.
By Proposition 2.9, we have
[…]
Note that the condition $\langle\pi^{r_{0}}D(I,\mathbf{r})B\sigma\rangle\leq_{s}\mathfrak{o}^{n}$ is equivalent to $\langle\pi^{r_{0}}D(I,\mathbf{r})B\rangle\leq_{s}\mathfrak{o}^{n}$, since $\sigma$ merely permutes the coordinates. Hence, using (2) and (4), we may rewrite the above as
$=\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\sum_{\sigma\in\mathcal{T}_{I,E}}\frac{\mu\bigl(\{B\in\operatorname{L}_{I}(\mathfrak{o}):\ \overline{B}=E\ \wedge\ \langle\pi^{r_{0}}D(I,\mathbf{r})B\rangle\leq_{s}\mathfrak{o}^{n}\}\bigr)}{\mu(\Gamma_{I,\mathbf{r}})}$
$=\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}\,q^{\sum_{\iota\in I}r_{\iota}\,\iota(n-\iota)}\,\mu\bigl(\{B\in\operatorname{L}_{I}(\mathfrak{o}):\ \overline{B}=E\ \wedge\ \langle\pi^{r_{0}}D(I,\mathbf{r})B\rangle\leq_{s}\mathfrak{o}^{n}\}\bigr).$
We now analyze the condition $\langle\pi^{r_{0}}D(I,\mathbf{r})B\rangle\leq_{s}\mathfrak{o}^{n}$. Write $D=D(I,\mathbf{r})$ and let $B_{1},\ldots,B_{n}$ denote the row vectors of $B$. Then the $i$-th row of $\pi^{r_{0}}DB$ is $\pi^{r_{0}}D_{ii}B_{i}$. Thus the submodule $\langle\pi^{r_{0}}DB\rangle$ is closed under multiplication if and only if, for all $1\leq i\leq i^{\prime}\leq n$, there exists a vector $v\in\mathfrak{o}^{n}$ such that
$\pi^{r_{0}}D_{ii}B_{i}\cdot\pi^{r_{0}}D_{i^{\prime}i^{\prime}}B_{i^{\prime}}=v\,\pi^{r_{0}}DB.$
The $j$-th entry of the vector $(B_{i}\cdot B_{i^{\prime}})B^{-1}$ is $\sum_{k=1}^{i}B_{ik}B_{i^{\prime}k}(B^{-1})_{kj}$. Consequently, the above condition is equivalent to
$\pi^{r_{0}}D_{ii}D_{i^{\prime}i^{\prime}}\left(\sum_{k\leq i}B_{ik}B_{i^{\prime}k}(B^{-1})_{kj}\right)\equiv 0\pmod{D_{jj}}$
for all $j=1,\ldots,n$. Equivalently, $\langle\pi^{r_{0}}DB\rangle$ is closed under multiplication if and only if
$\prod_{\begin{subarray}{c}\iota\in I\\ \iota\geq j\end{subarray}}\pi^{r_{\iota}}\;\Big|\;\pi^{r_{0}}\left(\prod_{\begin{subarray}{c}\iota\in I\\ \iota\geq i\end{subarray}}\pi^{r_{\iota}}\right)\left(\prod_{\begin{subarray}{c}\iota\in I\\ \iota\geq i^{\prime}\end{subarray}}\pi^{r_{\iota}}\right)\left(\sum_{k\leq i}B_{ik}B_{i^{\prime}k}(B^{-1})_{kj}\right)$
for all $1\leq i\leq i^{\prime}\leq n$ and all $j=1,\ldots,n$. Clearly, it suffices to verify this condition for $j<i$. ∎
Proceeding as in Example 2.10 and using Proposition 2.11, we obtain the following integral representation.
$\Xi_{n,I}\big(s_{0},(s_{\iota})_{\iota\in I}\big)$ $=\frac{1}{(1-q^{-1})^{|I|+1}}\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}\int\limits_{\begin{subarray}{c}(x_{0},(x_{\iota})_{\iota\in I},B)\in\mathscr{M}^{\prime}_{I}\\ \overline{B}=E\end{subarray}}|x_{0}|^{s_{0}n-1}\prod_{\iota\in I}|x_{\iota}|^{\iota(s_{\iota}-(n-\iota))-1}\,|d\mathbf{x}|\,|dB|.$
Here $\mathscr{M}^{\prime}_{I}\subseteq\mathfrak{o}\times\mathfrak{m}^{I}\times\operatorname{L}_{I}(\mathfrak{o})$ is the set of tuples $\big(x_{0},(x_{\iota})_{\iota\in I},B\big)$ satisfying the divisibility conditions
$\prod_{\begin{subarray}{c}\iota\in I\\ j\leq\iota<i\end{subarray}}x_{\iota}\;\Big|\;x_{0}\left(\prod_{\begin{subarray}{c}\iota\in I\\ i^{\prime}\leq\iota\end{subarray}}x_{\iota}\right)\mathscr{R}_{ii^{\prime}j}(B),\qquad\text{for all }(j,i,i^{\prime})\in[n]^{3}\text{ with }j<i\leq i^{\prime},$
where
$\mathscr{R}_{ii^{\prime}j}(B):=\sum_{k\leq i}B_{ik}B_{i^{\prime}k}(B^{-1})_{kj}.$
Finally, performing the change of variables $x_{\iota}\mapsto\pi x_{\iota}$ for $\iota\in I$, we obtain
$\Xi_{n,I}\big(s_{0},(s_{\iota})_{\iota\in I}\big)$ $=\frac{q^{-\sum_{\iota\in I}\iota(s_{\iota}-n+\iota)}}{(1-q^{-1})^{|I|+1}}\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}\int\limits_{\begin{subarray}{c}(x_{0},(x_{\iota})_{\iota\in I},B)\in\mathscr{M}_{I}\\ \overline{B}=E\end{subarray}}|x_{0}|^{s_{0}n-1}\prod_{\iota\in I}|x_{\iota}|^{\iota(s_{\iota}-(n-\iota))-1}\,|d\mathbf{x}|\,|dB|,$
where now $\mathscr{M}_{I}\subseteq\mathfrak{o}\times\mathfrak{o}^{I}\times\operatorname{L}_{I}(\mathfrak{o})$ is defined by
$\prod_{\begin{subarray}{c}\iota\in I\\ j\leq\iota<i\end{subarray}}\pi x_{\iota}\;\Big|\;x_{0}\left(\prod_{\begin{subarray}{c}\iota\in I\\ i^{\prime}\leq\iota\end{subarray}}\pi x_{\iota}\right)\mathscr{R}_{ii^{\prime}j}(B),\qquad\text{for all }(j,i,i^{\prime})\in[n]^{3}\text{ with }j<i\leq i^{\prime}.$
Remark 2.12.
In general, it seems difficult to give a closed formula for the numbers $\tau_{I,E}$. Note, however, that $\tau_{I,E}=\tau_{I,E^{\prime}}$ whenever $E$ and $E^{\prime}$ have the same pattern of zero entries. Thus, it suffices to formulate this problem over the field $\mathbb{F}_{2}$.
In the computations carried out in the subsequent sections, we will only need to evaluate certain sums of these quantities, which is often considerably simpler. For instance, we have
(8) $\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}=\#\mathcal{R}_{I}(\mathfrak{o}/\mathfrak{m})=\binom{n}{I}_{q}.$
3. The case #I=1\#I=1
Let $n\geq 2$ and $I=\{\iota\}$ for some $1\leq\iota<n$. In this section, we compute
$\Xi_{n,I}(s_{0},s_{\iota})=\frac{q^{-\iota(s_{\iota}-n+\iota)}}{(1-q^{-1})^{2}}\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}\int\limits_{\begin{subarray}{c}(x_{0},x_{\iota},B)\in\mathscr{M}_{I}\\ \overline{B}=E\end{subarray}}|x_{0}|^{s_{0}n-1}|x_{\iota}|^{\iota(s_{\iota}-(n-\iota))-1}|d\mathbf{x}||dB|$
and present some applications. In particular, we obtain new proofs for the formulas for $g_{n}(p^{n})$ and $g_{n}(p^{n+1})$ obtained in [11] and [1], and for the lower bound for the abscissa of convergence of $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ obtained in [7].
If $S$ is a subset of $\mathfrak{o}$, by $x|S$ we mean that $x|y$ for all $y\in S$.
We begin by describing $\mathscr{M}_{I}$. This is the set of those $(x_{0},x_{\iota},B)\in\mathfrak{o}^{2}\times\operatorname{L}_{I}(\mathfrak{o})$ satisfying the condition
$\pi x_{\iota}|x_{0}\mathscr{R}(B)$
where
$\mathscr{R}(B)$ $=\{\mathscr{R}_{i,i^{\prime},j}(B)\,:\,j\leq\iota<i\leq i^{\prime}\}$
$=\{B_{ij}(B_{ij}-1)\,:\,\iota+1\leq i\leq n,\ 1\leq j\leq\iota\}\cup\{B_{ij}B_{i^{\prime}j}\,:\,\iota+1\leq i<i^{\prime}\leq n,\ 1\leq j\leq\iota\}\cup\{0\}.$
We next compute a preliminary integral.
Definition 3.1.
A matrix is said to be very elemental if its coefficient are 0 or 1, and if in each column there is at most one 1.
Lemma 3.2.
Let $E=\begin{pmatrix}\operatorname{Id}_{\iota}&0\\ E^{\prime}&\operatorname{Id}_{n-\iota}\end{pmatrix}\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})$. Then
$\int\limits_{\begin{subarray}{c}\overline{B}=E\\ \pi x_{\iota}|\mathscr{R}(B)\end{subarray}}|x_{\iota}|^{s-1}|dx_{\iota}||dB|=\delta_{E}\cdot\frac{(1-q^{-1})q^{-(n-\iota)\iota}}{1-q^{-(s+(n-\iota)\iota)}}.$
where $\delta_{E}=1$ or 0 according to whether $E^{\prime}$ is very elemental or not.
Proof.
Let $B=\begin{pmatrix}\operatorname{Id}_{\iota}&0\\ \beta^{\prime}&\operatorname{Id}_{n-\iota}\end{pmatrix}\in\operatorname{L}_{I}(\mathfrak{o})$. The condition $\pi x_{\iota}|B_{ij}(B_{ij}-1)\,\forall i,j$ with $\iota+1\leq i\leq n$, $1\leq j\leq\iota$ translates as: each coefficient of $B^{\prime}$ is congruent to 0 or $1$ modulo $\pi x_{\iota}$. The condition $B_{ij}B_{i^{\prime}j}\,\forall i,i^{\prime},j$ with $\iota+1\leq i<i^{\prime}\leq n$, $1\leq j\leq\iota$ translates as: in each column of $B^{\prime}$ there is at most one unit. Hence, condition $\pi x_{\iota}|\mathscr{R}(\beta)$ is equivalent to: $B^{\prime}=\epsilon+\pi x_{\iota}B^{\prime\prime}$ for some very elementary matrix $\epsilon$ and some $B^{\prime\prime}\in\operatorname{Mat}_{(n-\iota)\times\iota}(\mathfrak{o})$. In particular, the integral is 0 if $\delta_{E}=0$. If $\delta_{E}=1$, then $\epsilon$ is completely determined, and we can apply the change of variables $B^{\prime}=\epsilon+\pi x_{\iota}B^{\prime\prime}$. We obtain
$\int\limits_{\begin{subarray}{c}\overline{\beta}=E\\ \pi x_{\iota}|\mathscr{R}(B)\end{subarray}}|x_{\iota}|^{s-1}|dx_{\iota}||dB|=\int|\pi x_{\iota}|^{(n-\iota)\iota}|x_{\iota}|^{s-1}|dx_{\iota}||dB^{\prime\prime}|=\frac{(1-q^{-1})q^{-(n-\iota)\iota}}{1-q^{-(s+(n-\iota)\iota)}}.$
Using our description of $\mathscr{M}_{I}$, we compute the next integral by partitioning the domain of integration:
$\int\limits_{\begin{subarray}{c}(x_{0},x_{\iota},B)\in\mathscr{M}_{I}\\ \overline{B}=E\end{subarray}}|x_{0}|^{s_{0}n-1}|x_{\iota}|^{\iota(s_{\iota}-(n-\iota))-1}|d\mathbf{x}||dB|=\left(\int\limits_{\begin{subarray}{c}\pi x_{\iota}|x_{0}\\ \overline{B}=E\end{subarray}}+\int\limits_{\begin{subarray}{c}x_{0}|x_{\iota}\\ \pi\frac{x_{\iota}}{x_{0}}|\mathscr{R}(B)\\ \overline{B}=E\end{subarray}}\right)|x_{0}|^{s_{0}n-1}|x_{\iota}|^{\iota(s_{\iota}-(n-\iota))-1}|d\mathbf{x}||dB|$
After applying the change of variables $(x_{0},x_{\iota})=(\pi x_{\iota}^{\prime}x_{0}^{\prime},x_{\iota}^{\prime})$ in the first integral and $(x_{0},x_{\iota})=(x_{0}^{\prime},x_{0}^{\prime}x_{\iota}^{\prime})$ in the second one, and renaming $x_{0}^{\prime}=x_{0}$ and $x_{\iota}^{\prime}=x_{\iota}$, the integral becomes:
$|\pi|^{s_{0}n}\int\limits_{\begin{subarray}{c}\overline{B}=E\end{subarray}}|x_{0}|^{s_{0}n-1}|x_{\iota}|^{s_{0}n+\iota(s_{\iota}-(n-\iota))-1}|d\mathbf{x}||dB|+\int\limits_{\begin{subarray}{c}\pi{x_{\iota}}|\mathscr{R}(B)\\ \overline{B}=E\end{subarray}}|x_{0}|^{s_{0}n+\iota(s_{\iota}-(n-\iota))-1}|x_{\iota}|^{\iota(s_{\iota}-(n-\iota))-1}|d\mathbf{x}||dB|.$
The first integral is straightforward, while for the second one, we use Lemma 3.2. The result is
$\frac{(1-q^{-1})^{2}q^{-(n-\iota)\iota}}{1-q^{-(s_{0}n+\iota(s_{\iota}-(n-\iota)))}}\left(\frac{q^{-s_{0}n}}{1-q^{-s_{0}n}}+\delta_{E}\frac{1}{1-q^{-\iota s_{\iota}}}\right)$
Next, we have to multiply this result by $\tau_{I,E}$ and compute the sum over $\mathcal{R}_{I}(\mathfrak{o}/\mathfrak{m})$. By (8), we have the formula
$\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}=\binom{n}{\iota}_{q}.$
The next lemma gives us the other necessary sum:
Lemma 3.3.
Let $1\leq\iota\leq n-1$ and $I=\{\iota\}$. Then
$\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\delta_{E}\,\tau_{I,E}=\sum_{\begin{subarray}{c}0\leq t_{1},\ldots,t_{n-\iota+1}\leq\iota\\ t_{1}+\cdots+t_{n-\iota+1}=\iota\end{subarray}}1^{t_{1}}\cdot 2^{t_{2}}\cdot\cdots\cdot(n-\iota+1)^{t_{n-\iota+1}},$
which is the coefficient of $z^{\iota}$ in the expansion of $\prod_{i=1}^{n-\iota+1}\frac{1}{1-iz}$.
Proof.
The sum to be computed is the number of $I$-reduced matrices over $\mathfrak{o}/\mathfrak{m}$ whose entries are only 0 or $1$ and such that there is at most two non-zero entries in each column. To count these matrices, we first choose the column positions $j_{\iota+1},\ldots,j_{n}$ for the leading 1’s of rows $\iota+1,\ldots,n$. The leading 1’s of the first $\iota$ rows are determined are then completely determined.
Next, we need to decide in which columns we are going to add an extra 1. Between columns $j_{\iota+1}$ and $j_{\iota+2}$ we can add extra 1’s without restriction in the $(\iota+1)$-th row, so there are $2^{j_{\iota+2}-j_{\iota+1}-1}$ possible choices. Between columns $j_{\iota+2}$ and $j_{\iota+3}$, we can only add extra 1’s in rows $\iota+1$ and $\iota+2$, so for each column between $j_{\iota+2}$ and $j_{\iota+3}$ we have 3 possibilities: to add no extra 1, to add an extra 1 in row $\iota+1$, or to add an extra 1 in row $\iota+2$. Hence, there are $3^{j_{\iota+3}-j_{\iota+2}-1}$ possible choices. Continuing the analysis in this way we find that
$\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\delta_{E}\,\tau_{I,E}$ $=\sum_{1\leq j_{\iota+1}<\cdots<j_{n}\leq n}2^{j_{\iota+2}-j_{\iota+1}-1}3^{j_{\iota+3}-j_{\iota+2}-1}\cdots(n-\iota)^{j_{n}-j_{n-1}-1}(n-\iota+1)^{n-j_{n}}$
$=\sum_{\begin{subarray}{c}0\leq t_{1},\ldots,t_{n-\iota+1}\leq\iota\\ t_{1}+\cdots+t_{n-\iota+1}=\iota\end{subarray}}1^{t_{1}}\cdot 2^{t_{2}}\cdot\cdots\cdot(n-\iota+1)^{t_{n-\iota+1}}.$
Putting everything together, we obtain:
Proposition 3.4.
Let $1\leq\iota\leq n-1$ and put $I=\{\iota\}$. Then
$\Xi_{n,I}(s_{0},s_{\iota})=\frac{q^{-\iota s_{\iota}}}{1-q^{-s_{0}n}q^{-\iota s_{\iota}}q^{\iota(n-\iota)}}\left(\binom{n}{\iota}_{q}\frac{q^{-s_{0}n}}{1-q^{-s_{0}n}}+\theta_{n,\iota}\frac{1}{1-q^{-\iota s_{\iota}}}\right)$
where $\theta_{n,\iota}$ is the coefficient of $z^{\iota}$ in the expansion of $\prod_{i=1}^{n-\iota+1}\frac{1}{1-iz}$.
3.1. Application I
We give a formula for
$\Xi_{2}(s)=\sum_{\Lambda\leq_{s}\mathfrak{o}^{2}}[\mathfrak{o}^{2}:\Lambda]^{-s}.$
Since the only non-empty $I\subseteq[2-1]$ is $I=\{1\}$, (6) gives
$\Xi_{2}(s)=\frac{1}{1-q^{-2s}}+\Xi_{2,\{1\}}(s,s).$
We now apply Proposition 3.4. Note that $\binom{2}{1}_{q}=1+q$ and the coefficient of $z$ in $\frac{1}{(1-z)(1-2z)}$ is $3$. Hence
$\Xi_{2}(s)$ $=\frac{1}{1-q^{-2s}}+\frac{q^{-s}}{1-q^{-3s+1}}\left((1+q)\frac{q^{-2s}}{1-q^{-2s}}+3\frac{1}{1-q^{-s}}\right)$
$=\frac{(1+q^{-s})^{2}}{(1-q^{-s})(1-q^{-3s+1})}.$
In particular, if $\mathfrak{o}=\mathbb{Z}_{p}$ we reobtain
$\zeta_{\mathbb{Z}_{p}^{3}}^{\scriptscriptstyle{1,<}}(s)=\zeta_{\mathbb{Z}_{p}^{2}}^{\scriptscriptstyle<}(s)=\frac{(1+p^{-s})^{2}}{(1-p^{-s})(1-p^{-3s+1})}\quad\text{and}\quad\zeta_{\mathbb{Z}^{3}}^{\scriptscriptstyle{1,<}}(s)=\frac{\zeta(s)^{3}\zeta(3s-1)}{\zeta(2s)^{2}}.$
3.2. Application II
We can also apply Proposition 3.4 to compute the first three non-zero coefficients of
$\sum_{\begin{subarray}{c}\Lambda\leq_{s}\mathfrak{o}^{n}\\ \Lambda\subseteq\pi\mathfrak{o}^{n}\end{subarray}}[\mathfrak{o}^{n}:\Lambda]^{-s}=b_{n}q^{-ns}+b_{n+1}q^{-(n+1)s}+b_{n+2}q^{-(n+2)s}+\cdots.$
The coefficient $b_{n}$ is 1, as $\pi\mathfrak{o}^{n}$ is the only submodule closed under multiplication of index $q^{n}$ included in $\pi\mathfrak{o}^{n}$.
To compute $b_{n+1}$, we note that a submodule $\Lambda\subset\mathfrak{o}^{n}$ included in $\pi\mathfrak{o}^{n}$ has index $q^{n+1}$ if and only if $\nu_{0}(\Lambda)=(\{1\},1,1)$. Hence, $b_{n+1}=h_{n}(\{1\},1,1)$, which is the coefficient of $q^{-ns_{0}}q^{-s_{1}}$ in the series $\Xi_{\{1\}}(s_{0},s_{1})$. By Proposition 3.4,
$b_{n}=\binom{n}{1}_{q}=\frac{1-q^{n}}{1-q}.$
To compute $b_{n+2}$ we first assume $n\geq 3$. We note that a submodule $\Lambda\subset\mathfrak{o}^{n}$ included in $\pi\mathfrak{o}^{n}$ has index $q^{n+2}$ if and only if $\nu(\Lambda)=(\{1\},1,2)$ or $\nu(\Lambda)=(\{2\},1,1)$. Hence, $b_{n+2}=h_{n}(\{1\},1,2)+h_{n}(\{2\},1,1)$. These terms are respectively the coefficients of $q^{-ns_{0}}q^{-2s_{1}}$ in the series $\Xi_{\{1\}}(s_{0},s_{1})$ and $\Xi_{\{2\}}(s_{0},s_{1})$. Hence, by Proposition 3.4,
$b_{n+2}=\theta_{n,1}q^{n-1}+\binom{n}{2}_{q}=\binom{n+1}{2}q^{n-1}+\binom{n}{2}_{q}.$
When $n=2$, $b_{n+2}=1+h_{n}(\{1\},1,2)$, the first term corresponding to the submodule $\pi^{2}\mathfrak{o}^{2}$. The above formula remains true as $\binom{n}{2}_{q}=1$ if $n=2$.
In particular, setting $\mathfrak{o}=\mathbb{Z}_{p}$, we obtain
$g_{n+1}(p^{n})=1,\quad g_{n+1}(p^{n+1})=\frac{1-q^{n}}{1-q},\quad g_{n+1}(p^{n+2})=\binom{n+1}{2}p^{n-1}+\binom{n}{2}_{q}.$
Remark 3.5.
Proposition 3.4 can be used to prove that for any $n$, any $\iota\in[n-1]$ and any $(r_{0},r)\in\mathbb{N}_{0}\times\mathbb{N}$, there exists a polynomial $P_{n,\iota,r_{0},r}(X)$ such that $h_{n}(\{\iota\},r_{0},r)=P_{n,\iota,r_{0},r}(q)$ for any compact discrete valuation ring $\mathfrak{o}$ with residue field of $q$ elements. More generally, we can ask if for any $n$, $I\subseteq[n-1]$ and $(r_{0},\mathbf{r})\in\mathbb{N}_{0}\times\mathbb{N}^{I}$, there exists a polynomial $P_{n,I,r_{0},\mathbf{r}}(X)$ such that $h_{n}(I,r_{0},\mathbf{r})=P_{n,I,r_{0},\mathbf{r}}(q)$ for any discrete valuation ring $\mathfrak{o}$ of residue field of $q$ elements. Certainly, a positive answer of this question would imply that for each $n\geq 1$ and each $e\geq 0$, there exists a polynomial $P_{n,e}(X)$ such that for any compact discrete valuation ring $\mathfrak{o}$ with residue field of $q$ elements, the coefficient of $q^{-es}$ in $\Xi_{n}(s)$ is $P_{n,e}(q)$.
3.3. Application III
We compute a lower bound for the abscissa of convergence of $\Xi_{n}(s)$. By Proposition 3.4, for each $\iota\in[n-1]$, the abscissa of convergence of $\Xi_{n,\{\iota\}}(s,s)$ is $\frac{\iota(n-\iota)}{n+\iota}$. Hence, $\Xi_{n}(s)$ has a pole at
$\max_{\iota\in[n-1]}\frac{\iota(n-\iota)}{\iota+n}.$
One can readily check that this number is $\geq\frac{n}{6}$ (take $\iota=\frac{n}{2}$ if $n$ is even or $\iota=\frac{n-1}{2}$ if $n$ is odd). Moreover, one can show that
$\lim\limits_{n\to\infty}\max_{\iota\in[n-1]}\frac{\iota(n-\iota)}{\iota+n}=3-2\sqrt{2}.$
Remark 3.6.
With a similar argument, or following the one given in [12, Theorem 4.1.3], one can show that if $L$ is the module $\mathfrak{o}^{n}$ with a $\mathfrak{o}$-bilinear multiplication, then the generating series of the submodules of $L$ that are closed under multiplication has abscissa of convergence bounded below by $\max_{\iota\in[n-1]}\frac{\iota(n-\iota)}{\iota+n}$.
4. The case #I=2\#I=2
Let $n\geq 3$ and let $I=\{i_{1},i_{2}\}$ with $1\leq i_{1}<i_{2}\leq n-1$. We consider the integral
$\Xi_{n,I}(s_{0},s_{1},s_{2})$ $=\frac{q^{-i_{1}(s_{1}-n+i_{1})-i_{2}(s_{2}-n+i_{2})}}{(1-q^{-1})^{3}}$
$\qquad\times\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}\int\limits_{\begin{subarray}{c}(x_{0},x_{1},x_{2},B)\in\mathscr{M}_{I}\\ \overline{B}=E\end{subarray}}|x_{0}|^{s_{0}n-1}|x_{1}|^{i_{1}(s_{1}-n+i_{1})-1}|x_{2}|^{i_{2}(s_{2}-n+i_{2})-1}|dx_{0}dx_{1}dx_{2}||dB|$
In §4.1, we describe the domain of integration and express the above integral as a finite sum of simpler integrals by partitioning the domain. These integrals are then computed explicitly in §4.2 in the special case $n=3$ and $I=\{1,2\}$. We finally use this calculation in 4.3 to write a formula for the cotype zeta function of $\mathfrak{o}^{4}$.
4.1. A first reduction
We begin by describing the domain $\mathscr{M}_{I}$. By definition, it consists of those
$(x_{0},x_{1},x_{2},B)\in\mathfrak{o}^{3}\times\operatorname{L}_{I}(\mathfrak{o})$
satisfying the following divisibility conditions:
$x_{1}$ $\mid x_{0}x_{2}\,\mathscr{R}_{ii^{\prime}j}(B)$ $\text{for }1\leq j\leq i_{1}<i\leq i^{\prime}\leq i_{2},$
$\pi x_{1}$ $\mid x_{0}\,\mathscr{R}_{ii^{\prime}j}(B)$ $\text{for }1\leq j\leq i_{1}<i\leq i_{2}<i^{\prime}\leq n,$
$\pi^{2}x_{1}x_{2}$ $\mid x_{0}\,\mathscr{R}_{ii^{\prime}j}(B)$ $\text{for }1\leq j\leq i_{1},\ i_{2}<i\leq i^{\prime}\leq n,$
$\pi x_{2}$ $\mid x_{0}\,\mathscr{R}_{ii^{\prime}j}(B)$ $\text{for }i_{1}<j\leq i_{2}<i\leq i^{\prime}\leq n.$
Writing
$B=\begin{pmatrix}\operatorname{Id}_{i_{1}}&0&0\\ (a_{ij})&\operatorname{Id}_{i_{2}-i_{1}}&0\\ (b_{ij})&(c_{ij})&\operatorname{Id}_{n-i_{2}}\end{pmatrix}$,
one easily checks that the following identities of sets hold.
$\{\mathscr{R}_{ii^{\prime}j}(B)\mid 1\leq j\leq i_{1}<i\leq i^{\prime}\leq i_{2}\}-\{0\}$ $=R_{1}(B),$
$\{\mathscr{R}_{ii^{\prime}j}(B)\mid 1\leq j\leq i_{1}<i\leq i_{2}<i^{\prime}\leq n\}-\{0\}$ $=R_{2}(B),$
$\{\mathscr{R}_{ii^{\prime}j}(B)\mid 1\leq j\leq i_{1},\ i_{2}<i\leq i^{\prime}\leq n\}-\{0\}$ $=R_{3}(B),$
$\{\mathscr{R}_{ii^{\prime}j}(B)\mid i_{1}<j\leq i_{2}<i\leq i^{\prime}\leq n\}-\{0\}$ $=R_{4}(B),$
where
$R_{1}(B):$ $=\{a_{ij}a_{i^{\prime}j}\mid 1\leq j\leq i_{1}<i<i^{\prime}\leq i_{2}\}\cup\{a_{ij}^{2}-a_{ij}\mid 1\leq j\leq i_{1}<i\leq i_{2}\}$
$R_{2}(B):$ $=\{a_{ij}(b_{i^{\prime}j}-c_{i^{\prime}i})\mid 1\leq j\leq i_{1}<i\leq i_{2}<i^{\prime}\leq n\},$
$R_{3}(B):$ $=\{b_{ij}b_{i^{\prime}j}-\sum_{k=i_{1}+1}^{i_{2}}a_{kj}c_{ik}c_{i^{\prime}k}\mid 1\leq j\leq i_{1},\ i_{2}<i<i^{\prime}\leq n\}$
$\quad\cup\{b_{ij}^{2}-b_{ij}-\sum_{k=i_{1}+1}^{i_{2}}a_{kj}(c_{ik}^{2}-c_{ik})\mid 1\leq j\leq i_{1},\ i_{2}<i\leq n\},$
$R_{4}(B):$ $=\{c_{ij}c_{i^{\prime}j}\mid i_{1}<j\leq i_{2}<i<i^{\prime}\leq n\}\cup\{c_{ij}^{2}-c_{ij}\mid i_{1}<j\leq i_{2}<i\leq n\}.$
For fixed $u_{0},u_{1},u_{2}\in\mathbb{C}$ and $E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})$ and a subset $\mathcal{C}\subset\mathfrak{o}^{3}\times\operatorname{L}_{I}(\mathfrak{o})$, set
$\mathcal{I}_{E}(\mathcal{C};\,u_{0},u_{1},u_{2}):=\frac{1}{(1-q^{-1})^{3}}\int\limits_{\begin{subarray}{c}(x_{0},x_{1},x_{2},B)\in\mathcal{C}\\ \overline{B}=E\end{subarray}}|x_{0}|^{u_{0}-1}\,|x_{1}|^{u_{1}-1}\,|x_{2}|^{u_{2}-1}\,|dx_{0}\,dx_{1}\,dx_{2}|\,|dB|.$
Thus
$\Xi_{n,I}(s_{0},s_{1},s_{2})=q^{-t_{1}-t_{2}}\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}\cdot\mathcal{I}_{E}(\mathcal{C}_{0};\,t_{0},t_{1},t_{2})$
where
$(t_{0},\,t_{1},\,t_{2})=\big(s_{0}n,\,i_{1}(s_{1}-n+i_{1}),\,i_{2}(s_{2}-n+i_{2})\big)$
and $\mathcal{C}_{0}$ is the set of $(x_{0},x_{1},x_{2},B)\in\mathfrak{o}^{3}\times\operatorname{L}_{I}(\mathfrak{o})$ such that
$x_{1}$ $\mid x_{0}x_{2}\,R_{1}(B),$ $\pi x_{1}$ $\mid x_{0}\,R_{2}(B),$
$\pi^{2}x_{1}x_{2}$ $\mid x_{0}\,R_{3}(B),$ $\pi x_{2}$ $\mid x_{0}\,R_{4}(B).$
Here $u\mid vS$ means $u\mid vs$ for all $s\in S$.
We compute $\mathcal{I}_{E}(\mathcal{C}_{0};\,t_{0},t_{1},t_{2})$ by successive partitions according to the relative divisibility of $x_{0},x_{1},x_{2}$. At each step we perform the corresponding change of variables and update the exponents.
Write
$\mathcal{C}_{0}=\mathcal{C}_{0,1}\sqcup\mathcal{C}_{0,2}\sqcup\mathcal{C}_{0,3}\sqcup\mathcal{C}_{0,4},$
where $\mathcal{C}_{0,1}$, $\mathcal{C}_{0,2}$, $\mathcal{C}_{0,3}$, $\mathcal{C}_{0,4}$ are defined respectively by the extra conditions:
  • $\pi^{2}x_{1}x_{2}\mid x_{0}$,
  • $\pi x_{1}|x_{0}$ and $\frac{x_{0}}{\pi x_{1}}\mid x_{2}$,
  • $x_{0}\mid x_{1}$ and $\pi x_{2}\mid x_{0}$,
  • $x_{0}\mid x_{1}$ and $x_{0}\mid x_{2}$.
On $\mathcal{C}_{0,1}$ we substitute $x_{0}=\pi^{2}x_{1}x_{2}x_{0}^{\prime}$ and obtain
$\mathcal{I}_{E}(\mathcal{C}_{0,1};t_{0},t_{1},t_{2})=|\pi|^{2t_{0}}\underbrace{\mathcal{I}_{E}\big(\mathfrak{o}^{3}\times\operatorname{L}_{I}(\mathfrak{o});t_{0},t_{0}+t_{1},t_{0}+t_{2}\big)}_{=:W_{1,E}}$
On $\mathcal{C}_{0,2}$ we substitute $x_{0}=\pi x_{1}x_{0}^{\prime}$ and $x_{2}=x_{0}^{\prime}x_{2}^{\prime}$, and obtain
$\mathcal{I}_{E}(\mathcal{C}_{0,2};t_{0},t_{1},t_{2})=|\pi|^{t_{0}}\mathcal{I}_{E}(\mathcal{C}_{2};t_{0}+t_{2},t_{0}+t_{1},t_{2})$
where $\mathcal{C}_{2}$ is defined by the conditions
$\pi x_{2}\mid R_{3}(B),$ $x_{2}\mid x_{1}\,R_{4}(B).$
We simplify this integral one step further. Write
$\mathcal{C}_{2}=\mathcal{C}_{2,1}\sqcup\mathcal{C}_{2,2},$
where $\mathcal{C}_{2,1}$ and $\mathcal{C}_{2,2}$ are defined respectively by the extra conditions
  • $x_{2}|x_{1}$, and
  • $\pi x_{1}|x_{2}$.
On $\mathcal{C}_{2,1}$ we substitute $x_{1}=x_{2}x_{1}^{\prime}$, and obtain
$\mathcal{I}_{E}(\mathcal{C}_{2,1}\,;t_{0}+t_{2},t_{0}+t_{1},t_{2})=\underbrace{\mathcal{I}_{E}\big(\pi x_{2}\mid R_{3}(B)\,;t_{0}+t_{2},t_{0}+t_{1},t_{0}+t_{1}+t_{2}\big)}_{=:W_{2,E}}$
On $\mathcal{C}_{2,2}$ we substitute $x_{2}=\pi x_{1}x_{2}^{\prime}$, and obtain
$\mathcal{I}_{E}(\mathcal{C}_{2,2}\,;t_{0}+t_{2},t_{0}+t_{1},t_{2})=|\pi|^{t_{2}}\underbrace{\mathcal{I}_{E}\left(\begin{array}[]{c}\pi^{2}x_{1}x_{2}\mid R_{3}(B)\\ \pi x_{2}\mid R_{4}(B)\end{array}\,;t_{0}+t_{2},t_{0}+t_{1}+t_{2},t_{2}\right)}_{=:W_{3,E}}$
On $\mathcal{C}_{0,3}$ we substitute $x_{0}=\pi x_{2}x_{0}^{\prime}$ and $x_{1}=\pi x_{2}x_{0}^{\prime}x_{1}^{\prime}$ and obtain
$\mathcal{I}_{E}(\mathcal{C}_{0,3}\,;t_{0},t_{1},t_{2})=|\pi|^{t_{0}+t_{1}}\mathcal{I}_{E}(\mathcal{C}_{3}\,;t_{0}+t_{1},\,t_{1},\,t_{0}+t_{1}+t_{2})$
where $\mathcal{C}_{3}$ is defined by the conditions
$x_{1}\mid x_{2}\,R_{1}(B),$ $\pi x_{1}\mid R_{2}(B),$ $\pi^{2}x_{1}x_{2}\mid R_{3}(B).$
We simplify this integral one step further. Write
$\mathcal{C}_{3}=\mathcal{C}_{3,1}\sqcup\mathcal{C}_{3,2},$
where $\mathcal{C}_{3,1}$ and $\mathcal{C}_{3,2}$ are defined respectively by the following conditions:
  • $x_{1}\mid x_{2}$,
  • $\pi x_{2}\mid x_{1}$.
On $\mathcal{C}_{3,1}$ we substitute $x_{2}=x_{1}x_{2}^{\prime}$ and obtain
$\mathcal{I}_{E}(\mathcal{C}_{3,1}\,;t_{0}+t_{1},\,t_{1},\,t_{0}+t_{1}+t_{2})=\underbrace{\mathcal{I}_{E}\left(\begin{array}[]{c}\pi x_{1}\mid R_{2}(B)\\ \pi^{2}x_{1}^{2}x_{2}\mid R_{3}(B)\end{array}\,;\,t_{0}+t_{1},\,t_{0}+2t_{1}+t_{2},\,t_{0}+t_{1}+t_{2}\right)}_{=:W_{4,E}}$
On $\mathcal{C}_{3,2}$ we substitute $x_{1}=\pi x_{2}x_{1}^{\prime}$ and obtain
$\mathcal{I}_{E}(\mathcal{C}_{3,2}\,;t_{0}+t_{1},\,t_{1},\,t_{0}+t_{1}+t_{2})=|\pi|^{t_{1}}\underbrace{\mathcal{I}_{E}\left(\begin{array}[]{c}\pi x_{1}\mid R_{1}(B)\\ \pi^{2}x_{1}x_{2}\mid R_{2}(B)\\ \pi^{3}x_{1}x_{2}^{2}\mid R_{3}(B)\end{array}\,;\,t_{0}+t_{1},\,t_{1},\,t_{0}+2t_{1}+t_{2}\right)}_{=:W_{5,E}}$
On $\mathcal{C}_{0,4}$ we substitute $x_{1}=x_{0}x_{1}^{\prime}$ and $x_{2}=x_{0}x_{2}^{\prime}$ and obtain
$\mathcal{I}_{E}(\mathcal{C}_{0,4}\,;\,t_{0},t_{1},t_{2})$ $=\mathcal{I}_{E}(\mathcal{C}_{4}\,;\,t_{0}+t_{1}+t_{2},\,t_{1},\,t_{2})$
where $\mathcal{C}_{4}$ is defined by the conditions
$x_{1}\mid x_{0}x_{2}R_{1}(B),$ $\pi x_{1}\mid R_{2}(B),$ $\pi^{2}x_{0}x_{1}x_{2}\mid R_{3}(B),$ $\pi x_{2}\mid R_{4}(B).$
We simplify this integral one step further. Write
$\mathcal{C}_{4}=\mathcal{C}_{4,1}\sqcup\mathcal{C}_{4,2}\sqcup\mathcal{C}_{4,3},$
where $\mathcal{C}_{4,1}$, $\mathcal{C}_{4,2}$, $\mathcal{C}_{4,3}$ are defined respectively by the following conditions
  • $x_{1}\mid x_{2}$,
  • $\pi x_{2}\mid x_{1}$ and $\frac{x_{1}}{x_{2}}\mid x_{0}$,
  • $\pi x_{0}x_{2}\mid x_{1}$.
On $\mathcal{C}_{4,1}$ we substitute $x_{2}=x_{1}x_{2}^{\prime}$ and obtain
$\mathcal{I}_{E}(\mathcal{C}_{4,1},\,;\,t_{0}+t_{1}+t_{2},\,t_{1},\,t_{2})=\underbrace{\mathcal{I}_{E}\left(\begin{array}[]{c}\pi x_{1}\mid R_{2}(B)\\ \pi^{2}x_{0}x_{1}^{2}x_{2}\mid R_{3}(B)\\ \pi x_{1}x_{2}\mid R_{4}(B)\end{array}\,;\,t_{0}+t_{1}+t_{2},\,t_{1}+t_{2},\,t_{2}\right)}_{=:W_{6,E}}.$
On $\mathcal{C}_{4,2}$ we substitute $x_{1}=\pi x_{2}x_{1}^{\prime}$, $x_{0}=\pi x_{1}^{\prime}x_{0}^{\prime}$ and obtain
$\mathcal{I}_{E}(\mathcal{C}_{4,2},\,;\,t_{0}+t_{1}+t_{2},\,t_{1},\,t_{2})=|\pi|^{t_{0}+2t_{1}+t_{2}}\underbrace{\mathcal{I}_{E}\left(\begin{array}[]{c}\pi^{2}x_{1}x_{2}\mid R_{2}(B)\\ \pi^{4}x_{0}x_{1}^{2}x_{2}^{2}\mid R_{3}(B)\\ \pi x_{2}\mid R_{4}(B)\end{array}\,;\,t_{0}+t_{1}+t_{2},\,t_{0}+2t_{1}+t_{2},\,t_{1}+t_{2}\right)}_{=:W_{7,E}}.$
On $\mathcal{C}_{4,3}$ we substitute $x_{1}=\pi x_{0}x_{2}x_{1}^{\prime}$ and obtain
$\mathcal{I}_{E}(\mathcal{C}_{4,3}\,;\,t_{0}+t_{1}+t_{2},\,t_{1},\,t_{2})=|\pi|^{t_{1}}\underbrace{\mathcal{I}_{E}\left(\begin{array}[]{c}\pi x_{1}\mid R_{1}(B)\\ \pi^{2}x_{0}x_{1}x_{2}\mid R_{2}(B)\\ \pi^{3}x_{0}^{2}x_{1}x_{2}^{2}\mid R_{3}(B)\\ \pi x_{2}\mid R_{4}(B)\end{array}\,;\,t_{0}+2t_{1}+t_{2},\,t_{1},\,t_{1}+t_{2}\right)}_{=:W_{8,E}}.$
Setting
$W_{i}:=\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}W_{i,E},$
and collecting the above identities, we obtain
$q^{t_{1}+t_{2}}\Xi_{I,\{i_{1},i_{2}\}}(s_{0},s_{1},s_{2})$ $=|\pi|^{2t_{0}}W_{1}+|\pi|^{t_{0}}W_{2}+|\pi|^{t_{0}+t_{2}}W_{3}+|\pi|^{t_{0}+t_{1}}W_{4}$
$\quad+|\pi|^{t_{0}+2t_{1}}W_{5}+W_{6}+|\pi|^{t_{0}+2t_{1}+t_{2}}W_{7}+|\pi|^{t_{1}}W_{8}.$
(Recall that $(t_{0},t_{1},t_{2})=\big(s_{0}n,\,i_{1}(s_{1}-n+i_{1}),\,i_{2}(s_{2}-n+i_{2})\big)$.)
Note that
$W_{1}$ $=\sum_{E\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})}\tau_{I,E}\int\limits_{\mathfrak{o}^{3}}|x_{0}|^{t_{0}-1}|x_{1}|^{t_{0}+t_{1}-1}|x_{2}|^{t_{0}+t_{2}-1}|dx_{0}dx_{1}dx_{2}|\cdot\int\limits_{\overline{B}=E}|dB|$
(9) $=\binom{n}{i_{2}}_{q}\binom{i_{2}}{i_{1}}_{q}\frac{q^{-(i_{1}(i_{2}-i_{1})+i_{2}(n-i_{2}))}}{(1-q^{-t_{0}})(1-q^{-(t_{0}+t_{1})})(1-q^{-(t_{0}+t_{2})})}.$
Our goal in the next section will be to compute $W_{2},\ldots,W_{8}$ when $n=3$ and $I=\{1,2\}$, and obtain $\Xi_{3,\{1,2\}}(s_{0},s_{1},s_{2})$.
4.2. The case $n=3$ and $I=\{1,2\}$
In this section we assume that $n=3$ and $I=\{1,2\}$. We are going to compute $W_{i}=W_{i}(t_{0},t_{1},t_{2})$ explicitly for $i=1,\ldots,8$, and then we will express $\Xi_{3,\{1,2\}}(s_{0},s_{1},s_{2})$ as a rational function in
$q,\qquad X=q^{-t_{0}},\qquad Y=q^{-t_{1}},\qquad Z=q^{-t_{2}},$
where $(t_{0},t_{1},t_{2})=(3s_{0},s_{1}-2,2s_{2}-2)$.
The computation of $W_{1}$ is a particular case of (9), so we focus on the computation of $W_{i}$ for $i=2,\ldots,8$. To begin with, we fix
$E=\begin{pmatrix}1&0&0\\ \alpha&1&0\\ \beta&\gamma&1\end{pmatrix}\in\operatorname{L}_{I}(\mathfrak{o}/\mathfrak{m})$
and consider the computation of the integrals $W_{i,E}$ defined in the previous section.
It is easy to check that for a matrix
$B=\begin{pmatrix}1&0&0\\ a&1&0\\ b&c&1\end{pmatrix}$
we have
$R_{1}(B)=\{a^{2}-a\},\qquad R_{2}(B)=\{a(b-c)\},\qquad R_{3}(B)=\{b^{2}-b-a(c^{2}-c)\},\qquad R_{4}(B)=\{c^{2}-c\}.$
Thus, for the computation of $W_{i,E}$ for all $i=2,\ldots,8$ we need to consider the computation of the integral
$S_{\alpha,\beta,\gamma}(u_{1},u_{2},u_{3},u_{4}):=\int\limits_{\begin{subarray}{c}\overline{a}=\alpha,\ \overline{b}=\beta,\ \overline{c}=\gamma\\ u_{1}\mid a^{2}-a\\ u_{2}\mid a(b-c)\\ u_{3}\mid b^{2}-b-a(c^{2}-c)\\ u_{4}\mid c^{2}-c\end{subarray}}|d(a,b,c)|$
for certain expressions $u_{1},u_{2},u_{3},u_{4}\in\mathfrak{o}$ on the variables $x_{0},x_{1},x_{2}$. To compute this integral we do some preparation.
For $\epsilon\in\mathfrak{o}/\mathfrak{m}$ we define
$\chi_{0}(\epsilon)=\begin{cases}1,&\text{if }\epsilon=0,\\ 0,&\text{if }\epsilon\neq 0,\end{cases}\qquad\text{and}\qquad\chi^{\times}(\epsilon)=1-\chi_{0}(\epsilon).$
In other words, $\chi_{0}$ is the characteristic function of the zero element of $\mathfrak{o}/\mathfrak{m}$ and $\chi^{\times}$ is that of the units. We extend the definition of $\chi_{0}$ to more variables by setting
$\chi_{0}(\epsilon_{1},\ldots,\epsilon_{r})=\chi_{0}(\epsilon_{1})\cdots\chi_{0}(\epsilon_{r}).$
We record two important observations.
  1. ((a))
    For any $c\in\mathfrak{o}$,
    $c^{2}-c=\big(c-\chi_{0}(\overline{c})\big)\big(c-\chi^{\times}(\overline{c})\big).$
    Moreover, $c-\chi_{0}(\overline{c})$ is always a unit. Thus, given $u\in\mathfrak{o}$,
    $u\mid c^{2}-c\Longleftrightarrow u\mid c-\chi^{\times}(\overline{c}).$
  2. ((b))
    Setting
    $\alpha_{1}:=\alpha\frac{\gamma-\chi_{0}(\gamma)}{\beta-\chi_{0}(\beta)},\qquad\beta_{1}:=\beta-\chi^{\times}(\beta),\qquad\gamma_{1}:=\gamma-\chi^{\times}(\gamma),$
    the map
    $(a,b,c)\longmapsto\left(a\frac{c-\chi_{0}(\gamma)}{b-\chi_{0}(\beta)},\ b-\chi^{\times}(\beta),\ c-\chi^{\times}(\gamma)\right)$
    is a bijection
    $\{(a,b,c)\in\mathfrak{o}^{3}:\ \overline{a}=\alpha,\ \overline{b}=\beta,\ \overline{c}=\gamma\}\cong\{(a_{1},b_{1},c_{1})\in\mathfrak{o}^{3}:\ \overline{a_{1}}=\alpha_{1},\ \overline{b_{1}}=\beta_{1},\ \overline{c_{1}}=\gamma_{1}\}.$
    Moreover, it is a diffeomorphism whose Jacobian has determinant $\frac{c-\chi_{0}(\gamma)}{b-\chi_{0}(\beta)}$, which is a unit.
Lemma 4.1.
Let $u_{1},u_{2},u_{3},u_{4}\in\mathfrak{o}$ satisfy
$\pi\mid u_{3},\qquad u_{2}u_{4}\mid u_{3}.$
Assume in addition that if $\pi\mid u_{1}$ then $u_{1}u_{4}\mid u_{2}$. Then for any $\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}$, the integral
$S=S_{\alpha,\beta,\gamma}(u_{1},u_{2},u_{3},u_{4})$
is given by
[…]
Proof.
Since $\pi|u_{3}$, the integral will be 0 unless
$\beta^{2}-\beta-\alpha(\gamma^{2}-\gamma)=0,$
or equivalently, $\beta_{1}-\alpha_{1}\gamma_{1}=0$. So we will assume this condition.
We claim that in each of the cases of the lemma, the correspondence described in (b) above induces a bijection
$\left\{(a,b,c)\in\mathfrak{o}^{3}\,:\,\begin{matrix}\overline{a}=\alpha,\ \overline{b}=\beta,\ \overline{c}=\gamma\\ u_{1}\mid a^{2}-a\\ u_{2}\mid a(b-c)\\ u_{3}\mid b^{2}-b-a(c^{2}-c)\\ u_{4}\mid c^{2}-c\end{matrix}\right\}\cong\left\{(a_{1},b_{1},c_{1})\in\mathfrak{o}^{3}\,:\,\begin{matrix}\overline{a}_{1}=\alpha_{1},\ \overline{c}_{1}=\gamma_{1}\\ u_{1}\mid a_{1}-\chi^{\times}(\alpha)\\ u_{2}\mid a_{1}\bigl(b_{1}-c_{1}+\chi^{\times}(\beta)-\chi^{\times}(\gamma)\bigr)\\ u_{3}\mid b_{1}-a_{1}c_{1}\\ u_{4}\mid c_{1}\end{matrix}\right\}.$
To prove this claim, it suffices to explain why the condition $\overline{b}_{1}=\beta_{1}$ is omitted on the right-hand side and why $u_{1}\mid a^{2}-a$, which is equivalent to $a-\chi^{\times}(\alpha)$, translates to $u_{1}\mid a_{1}-\chi^{\times}(\alpha)$.
The condition $\overline{b}_{1}=\beta_{1}$ follows from $\beta_{1}=\alpha_{1}\gamma_{1}$, $\pi\mid u_{3}\mid b_{1}-a_{1}c_{1}$, and $(\overline{a}_{1},\overline{c}_{1})=(\alpha_{1},\gamma_{1})$.
As for the translation of $u_{1}\mid a-\chi^{\times}(\alpha)$, we only need to consider those cases where $\pi\mid u_{1}$. Note that in those cases we also have $u_{1}\mid u_{2}$, and in particular $\pi\mid u_{2}$. When $\alpha=0$ then $\chi^{\times}(\alpha)=0$ and $u_{1}\mid a$ is equivalent to $u_{1}\mid a_{1}$. If $\alpha\neq 0$, then looking at the conditions in both sets, we find that they are empty unless $\beta=\gamma$. Suppose that this is the case. Then the conditions $u_{1}\mid a-1$ and $u_{2}\mid a(b-c)$ translate to
$u_{1}\mid a_{1}\frac{b_{1}+\chi^{\times}(\beta)-\chi_{0}(\beta)}{c_{1}+\chi^{\times}(\gamma)-\chi_{0}(\gamma)}-\chi^{\times}(\alpha)\qquad\text{and}\qquad u_{2}\mid b_{1}-c_{1}.$
Since $u_{1}\mid u_{2}$ and $\beta=\gamma$, we can use the second condition to reduce the first condition to $u_{1}\mid a_{1}-\chi^{\times}(\alpha)$. This completes the proof of the claim.
We can now apply a change of variables to compute $S$. Note that we may add the condition $u_{4}\mid b_{1}$ on the right-hand side since it follows from the last two conditions and the hypothesis $u_{4}\mid u_{3}$. Hence (always assuming $\beta^{2}-\beta-\alpha(\gamma^{2}-\gamma)=0$, otherwise the integral is zero),
$S$ $=\int\limits_{\begin{subarray}{c}\overline{a}=\alpha_{1},\ \overline{c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ u_{2}\mid a(b-c+\chi^{\times}(\beta)-\chi^{\times}(\gamma))\\ u_{3}\mid b-ac\\ u_{4}\mid c,\ u_{4}\mid b\end{subarray}}|d(a,b,c)|=\int\limits_{\begin{subarray}{c}\overline{a}=\alpha_{1},\ \overline{u_{4}c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ u_{2}\mid a(u_{4}b-u_{4}c+\chi^{\times}(\beta)-\chi^{\times}(\gamma))\\ u_{3}\mid u_{4}b-au_{4}c\end{subarray}}|u_{4}|^{2}\,|d(a,b,c)|$
$=\int\limits_{\begin{subarray}{c}\overline{a}=\alpha_{1},\ \overline{u_{4}c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ u_{2}\mid a((a-1)u_{4}c+\chi^{\times}(\beta)-\chi^{\times}(\gamma))\\ \frac{u_{3}}{u_{4}}\mid b-ac\end{subarray}}|u_{4}|^{2}\,|d(a,b,c)|=\int\limits_{\begin{subarray}{c}\overline{a}=\alpha_{1},\ \overline{u_{4}c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ u_{2}\mid a((a-1)u_{4}c+\chi^{\times}(\beta)-\chi^{\times}(\gamma))\end{subarray}}|u_{4}|\,|u_{3}|\,|d(a,c)|.$
  • Assume that $u_{2}$ is a unit. Then $u_{1}$ is a unit and the result is
    $S=\int\limits_{\overline{a}=\alpha_{1},\,\overline{u_{4}c}=\gamma_{1}}|u_{3}u_{4}||d(a,c)|=|\pi u_{3}u_{4}|\mu(\{c\,:\,\overline{u_{4}c}=\gamma_{1}\})=\begin{cases}|\pi|^{2}|u_{3}u_{4}|&\text{if $u_{4}$ is a unit,}\\ \chi_{0}(\gamma^{2}-\gamma)|\pi||u_{3}u_{4}|&\text{if $\pi\mid u_{4}$.}\end{cases}$
  • Assume that $\pi\mid u_{2}$. Then
    $S$ $=\int\limits_{\begin{subarray}{c}\overline{a}=\alpha_{1},\ \overline{u_{4}c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ u_{2}\mid a\end{subarray}}|u_{4}|\,|u_{3}|\,|d(a,c)|+\!\!\int\limits_{\begin{subarray}{c}\overline{a}=\alpha_{1},\ \overline{u_{4}c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ \pi a\mid u_{2}\\ u_{2}\mid a((a-1)u_{4}c+\chi^{\times}(\beta)-\chi^{\times}(\gamma))\end{subarray}}|u_{4}|\,|u_{3}|\,|d(a,c)|$
    $=:S_{1}+S_{2}.$
    We compute $S_{1}$. Since $\pi|u_{2}$, we have $S_{1}=0$ unless $\alpha_{1}=0$, or equivalently, $\alpha=0$. In that case, the conditions $\overline{a}=\alpha_{1}$ and $u_{1}\mid a$ follow from $u_{2}\mid a$ since $u_{1}\mid u_{2}$ and $\pi\mid u_{2}$, so they can be omitted. Hence
    $S_{1}$ $=\chi_{0}(\alpha)\int\limits_{\begin{subarray}{c}\overline{u_{4}c}=\gamma_{1}\\ u_{2}|a\end{subarray}}|u_{3}u_{4}||d(a,c)|=\chi_{0}(\alpha)|u_{2}u_{3}u_{4}|\,\cdot\,\int\limits_{\overline{u_{4}c}=\gamma_{1}}|dc|$
    $=\begin{cases}\chi_{0}(\alpha)|\pi||u_{2}u_{3}u_{4}|&\text{if $u_{4}$ is a unit,}\\ \chi_{0}(\alpha,\gamma^{2}-\gamma)|u_{2}u_{3}u_{4}|&\text{if $\pi\mid u_{4}$.}\end{cases}$
    We now compute $S_{2}$. We claim that the integral is zero unless $\beta=\gamma$. Indeed, the first, third, and fourth conditions defining the domain imply
    $(\alpha_{1}-1)\gamma_{1}+\chi^{\times}(\beta)-\chi^{\times}(\gamma)=0.$
    Since $\beta_{1}=\alpha_{1}\gamma_{1}$, this can be rewritten as
    $\beta_{1}-\gamma_{1}+\chi_{0}(\beta_{1})-\chi_{0}(\gamma_{1})=0,$
    i.e. $\beta-\gamma=0$.
    Thus, for the computation we assume $\beta=\gamma$, which implies $\alpha_{1}=\alpha$. The last condition in the domain of integration can then be written as
    $u_{2}\mid\big(a-\chi^{\times}(\alpha)\big)u_{4}c.$
    To proceed with the calculation of $S_{2}$ and $S$, we analyze the cases $u_{2}\mid u_{4}$ and $\pi u_{4}\mid u_{2}$ separately.
    • Assume that $u_{2}\mid u_{4}$. Then $\pi\mid u_{4}$ and $u_{1}$ is a unit. Note that the condition $\overline{u_{4}c}=\gamma_{1}$ implies that $S_{2}=0$ unless $\gamma_{1}=0$, and in the latter case the condition is redundant and can be omitted. Therefore,
      $S_{2}=\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)\int\limits_{\begin{subarray}{c}\overline{a}=\alpha\\ \pi a\mid u_{2}\end{subarray}}|u_{3}u_{4}||d(a,c)|.$
      If $\alpha=0$, we may substitute $a=\pi a^{\prime}$ and obtain
      $S_{2}=\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)|\pi|\,|u_{3}u_{4}|\int_{\pi^{2}a\mid u_{2}}|da|=\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)|\pi|\,|u_{3}u_{4}|\bigl(1-|u_{2}/\pi|\bigr).$
      If $\alpha\neq 0$, then the integral is defined only by $\overline{a}=\alpha$, and so $S_{2}=|\pi|\,|u_{3}u_{4}|$. Thus,
      $S_{2}$ $=\chi_{0}(\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|\pi|\,|u_{3}u_{4}|\bigl(1-|u_{2}/\pi|\bigr)+\chi^{\times}(\alpha)\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)\,|\pi|\,|u_{3}u_{4}|$
      $=\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)|\pi u_{3}u_{4}|-\chi_{0}(\alpha,\beta-\gamma,\gamma^{2}-\gamma)|u_{2}u_{3}u_{4}|.$
      Taking into account that $S=\chi_{0}\big(\beta^{2}-\beta-\alpha(\gamma^{2}-\gamma)\big)S$, we obtain that if $\pi\mid u_{2}\mid u_{4}$, then for any $\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}$ we have
      $S$ $=\chi_{0}(\alpha,\beta^{2}-\beta,\gamma^{2}-\gamma)|u_{2}u_{3}u_{4}|+\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)|\pi u_{3}u_{4}|-\chi_{0}(\alpha,\beta-\gamma,\gamma^{2}-\gamma)|u_{2}u_{3}u_{4}|$
      $=\big(\chi_{0}(\alpha,\beta,\gamma-1)+\chi_{0}(\alpha,\beta-1,\gamma)\big)|u_{2}u_{3}u_{4}|+\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)|\pi u_{3}u_{4}|$
    • We now assume that $\pi u_{4}\mid u_{2}$. Then
      $S_{2}$ $=\chi_{0}(\beta-\gamma)\int\limits_{\begin{subarray}{c}\overline{a}=\alpha,\ \overline{u_{4}c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ \pi a\mid u_{2}\\ u_{2}\mid(a-\chi^{\times}(\alpha))u_{4}c\end{subarray}}|u_{4}|\,|u_{3}|\,|d(a,c)|$
      $=\chi_{0}(\beta-\gamma)\underbrace{\int\limits_{\begin{subarray}{c}\overline{a}=\alpha,\ \overline{u_{4}c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ \pi a\mid u_{2}\\ u_{2}\mid(a-\chi^{\times}(\alpha))u_{4}\end{subarray}}|u_{4}|\,|u_{3}|\,|d(a,c)|}_{=:S_{2,1}}+\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)\underbrace{\int\limits_{\begin{subarray}{c}\overline{a}=\alpha,\ \overline{u_{4}c}=\gamma_{1}\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ \pi a\mid u_{2}\\ \pi(a-\chi^{\times}(\alpha))u_{4}\mid u_{2}\\ \frac{u_{2}}{(a-\chi^{\times}(\alpha))u_{4}}\mid c\end{subarray}}|u_{4}|\,|u_{3}|\,|d(a,c)|}_{=:S_{2,2}}.$
      We compute $S_{2,1}$. We show first that the condition $u_{1}\mid a-\chi^{\times}(\alpha)$ can be removed. This is obvious if $u_{1}$ is a unit. If $\pi\mid u_{1}$, then this condition follows from the last condition and the hypothesis $u_{1}u_{4}\mid u_{2}$. Finally, conditions $\pi u_{4}\mid u_{2}$ and $u_{2}\mid\big(a-\chi^{\times}(\alpha)\big)u_{4}$ imply $\alpha-\chi^{\times}(\alpha)=0$, hence $\alpha=0$ or $1$, so the condition $\overline{a}=\alpha$ can be omitted. Therefore,
      $S_{2,1}$ $=\chi_{0}(\alpha^{2}-\alpha)\int\limits_{\begin{subarray}{c}\pi a\mid u_{2}\\ \frac{u_{2}}{u_{4}}\mid a-\chi^{\times}(\alpha)\end{subarray}}|u_{3}|\,|u_{4}|\,|da|\,\cdot\,\int\limits_{\overline{u_{4}c}=\gamma_{1}}|dc|$
      $=\chi_{0}(\alpha^{2}-\alpha)\,\int\limits_{\begin{subarray}{c}\pi\bigl(a\frac{u_{2}}{u_{4}}+\chi^{\times}(\alpha)\bigr)\mid u_{2}\end{subarray}}|u_{2}u_{3}||da|\,\cdot\,\int\limits_{\overline{u_{4}c}=\gamma_{1}}|dc|$
      $=\left(\chi_{0}(\alpha)\int_{\pi a\mid u_{4}}|u_{2}u_{3}||da|+\chi_{0}(\alpha-1)\int|u_{2}u_{3}||da|\right)\cdot\int_{\overline{u_{4}c}=\gamma_{1}}|dc|$
      $=\big(\chi_{0}(\alpha)|u_{2}u_{3}|(1-|u_{4}|)+\chi_{0}(\alpha-1)|u_{2}u_{3}|\big)\cdot\int_{\overline{u_{4}c}=\gamma_{1}}|dc|.$
      We now compute $S_{2,2}$. The condition $\overline{u_{4}c}=\gamma_{1}$ and the last condition in the domain imply $\gamma_{1}=0$, that is $\gamma^{2}-\gamma=0$, in which case $\overline{u_{4}c}=\gamma_{1}$ can be omitted. The condition $\pi a\mid u_{2}$ can also be omitted: this is clear if $\alpha\neq 0$, while if $\alpha=0$ then $\chi^{\times}(\alpha)=0$, hence it follows from $\pi au_{4}\mid u_{2}$. After integrating with respect to $c$, we obtain
      $S_{2,2}$ $=\chi_{0}(\gamma^{2}-\gamma)\,|u_{2}u_{3}|\int\limits_{\begin{subarray}{c}\overline{a}=\alpha\\ u_{1}\mid a-\chi^{\times}(\alpha)\\ a-\chi^{\times}(\alpha)\mid\frac{u_{2}}{\pi u_{4}}\end{subarray}}\frac{1}{|a-\chi^{\times}(\alpha)|}\,|da|$
      $=\chi_{0}(\gamma^{2}-\gamma)\,|u_{2}u_{3}|\int\limits_{\begin{subarray}{c}\overline{a}=\alpha-\chi^{\times}(\alpha)\\ u_{1}\mid a\mid\frac{u_{2}}{\pi u_{4}}\end{subarray}}\frac{1}{|a|}\,|da|.$
      If $\alpha=0$ or $1$, then the last integral equals
      $\int\limits_{\begin{subarray}{c}\overline{a}=0\\ u_{1}\mid a\mid\frac{u_{2}}{\pi u_{4}}\end{subarray}}\frac{1}{|a|}\,|da|=\int\limits_{\pi^{1-\chi_{0}(\overline{u}_{1})}u_{1}\mid a\mid\frac{u_{2}}{\pi u_{4}}}\frac{1}{|a|}\,|da|=\int\limits_{\pi a\mid\frac{u_{2}}{\pi^{1-\chi_{0}(\overline{u}_{1})}u_{1}u_{4}}}\frac{1}{|a|}\,|da|.$
      If $\alpha\neq 0,1$, then $u_{1}$ is necessarily a unit (otherwise the integral is 0), and in this case the integral is $|\pi|$. Summarizing,
      $S_{2,2}=\chi_{0}(\alpha^{2}-\alpha,\gamma^{2}-\gamma)\,|u_{2}u_{3}|\int_{\pi a\mid\frac{u_{2}}{\pi^{1-\chi_{0}(\overline{u}_{1})}u_{1}u_{4}}}\frac{1}{|a|}\,|da|+\chi^{\times}(\overline{u}_{1})\chi^{\times}(\alpha^{2}-\alpha)\chi_{0}(\gamma^{2}-\gamma)\,|\pi|\,|u_{2}u_{3}|.$
      Finally, using that $S=\chi_{0}\big(\beta^{2}-\beta-\alpha(\gamma^{2}-\gamma)\big)S$ and the formulas for $S_{1}$, $S_{2,1}$ and $S_{2,2}$, we obtain that if $\pi u_{4}\mid u_{2}$, then for any $\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}$ we have
        If $u_{4}$ is a unit then
        $S$ $=\left(\chi_{0}(\alpha,\beta^{2}-\beta)+\chi_{0}(\alpha-1,\beta-\gamma)+\chi^{\times}(\overline{u}_{1})\chi^{\times}(\alpha^{2}-\alpha)\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)\right)|u_{2}u_{3}||\pi|$
        $\quad+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|u_{2}u_{3}|\int_{\pi a\mid\frac{u_{2}}{\pi^{1-\chi_{0}(\overline{u}_{1})}u_{1}}}\frac{1}{|a|}\,|da|.$
        If $\pi\mid u_{4}$ then
        $S$ $=\left(\chi_{0}(\alpha,\beta^{2}-\beta,\gamma^{2}-\gamma)-\chi_{0}(\alpha,\beta-\gamma,\gamma^{2}-\gamma)\right)|u_{2}u_{3}u_{4}|+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)|u_{2}u_{3}|$
        $\quad+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|u_{2}u_{3}|\int_{\pi a\mid\frac{u_{2}}{\pi^{1-\chi_{0}(\overline{u}_{1})}u_{1}u_{4}}}\frac{1}{|a|}\,|da|$
        $\quad+\chi^{\times}(\overline{u}_{1})\chi^{\times}(\alpha^{2}-\alpha)\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)\,|\pi|\,|u_{2}u_{3}|.$
      The proof of the lemma follows clearly from these computations.
For $\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}$ we write $\tau_{\alpha,\beta,\gamma}$ for $\tau_{I,E}$, where
$E=\begin{pmatrix}1&0&0\\ \alpha&1&0\\ \beta&\gamma&1\end{pmatrix}$.
In other words, $\tau_{\alpha,\beta,\gamma}$ is the number of $\{1,2\}$-reduced matrices in $\operatorname{GL}_{3}(\mathfrak{o}/\mathfrak{m})$ obtained from $E$ by permutation of columns. To compute $\tau_{\alpha,\beta,\gamma}$ explicitly, we list all possible $\{1,2\}$-reduced matrices together with their associated lower triangular matrices.
  1. ((a))
    $\begin{pmatrix}0&0&1\\ 0&1&*\\ 1&*&*\end{pmatrix}\;\Rightarrow\;\begin{pmatrix}1&0&0\\ *&1&0\\ *&*&1\end{pmatrix}$
  2. ((b))
    $\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&*&*\end{pmatrix}\;\Rightarrow\;\begin{pmatrix}1&0&0\\ 0&1&0\\ *&*&1\end{pmatrix}$
  3. ((c))
    $\begin{pmatrix}0&0&1\\ 1&0&*\\ 0&1&*\end{pmatrix}\;\Rightarrow\;\begin{pmatrix}1&0&0\\ *&1&0\\ *&0&1\end{pmatrix}$
  4. ((d))
    $\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&*\end{pmatrix}\;\Rightarrow\;\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&*&1\end{pmatrix}$
  5. ((e))
    $\begin{pmatrix}0&1&0\\ 1&*&0\\ 0&0&1\end{pmatrix}\;\Rightarrow\;\begin{pmatrix}1&0&0\\ *&1&0\\ 0&0&1\end{pmatrix}$
  6. ((f))
    $\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\;\Rightarrow\;\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$
From this list we deduce that
[…]
Let
$S_{i}(x_{0},x_{1},x_{2})=\sum_{\alpha,\beta,\gamma}\tau_{\alpha,\beta,\gamma}S_{\alpha,\beta,\gamma}(u_{1},u_{2},u_{3},u_{4})$
where $(u_{1},u_{2},u_{3},u_{4})$ is
  • $(1,1,\pi x_{2},)$ if $i=2$,
  • $(1,1,\pi^{2}x_{1}x_{2},\pi x_{2})$ if $i=3$,
  • $(1,\pi x_{1},\pi^{2}x_{1}^{2}x_{1},1)$ if $i=4$,
  • $(\pi x_{1},\pi^{2}x_{1}x_{2},\pi^{3}x_{1}x_{2}^{2},1)$ if $i=5$,
  • $(1,\pi x_{1},\pi^{2}x_{0}x_{1}^{2}x_{2},\pi x_{1})$ if $i=6$,
  • $(1,\pi^{2}x_{1}x_{2},\pi^{4}x_{0}x_{1}^{2}x_{2}^{2},\pi x_{2})$ if $i=7$,
  • $(\pi x_{1},\pi^{2}x_{0}x_{1}x_{2},\pi^{3}x_{0}^{2}x_{1}x_{2}^{2},\pi x_{2})$ if $i=8$.
Let
$\mathcal{W}_{i}(r_{0},r_{1},r_{2})=\frac{1}{(1-q^{-1})^{3}}\int|x_{0}|^{r_{1}-1}|x_{1}|^{r_{2}-1}|x_{2}|^{r_{3}-1}S_{i}(x_{0},x_{1},x_{2})|d(x_{0},x_{2},x_{2})|.$
Then
$W_{i}=W_{i}(t_{0},t_{1},t_{2})=\mathcal{W}_{i}(r_{0},r_{1},r_{2})$
for some linear functions $r_{0},r_{1},r_{2}$ on $t_{0},t_{1},t_{2}$.
To compute $S_{i}(x_{0},x_{1},x_{2})$ we only need to apply Lemma 4.1 and the following computation of weighted coefficients, whose verification is straightforward.
  • $\sum_{\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}}\tau_{\alpha,\beta,\gamma}\,\chi_{0}\!\bigl(\beta^{2}-\beta-\alpha(\gamma^{2}-\gamma)\bigr)=q^{2}+8q+1$.
  • $\sum_{\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}}\tau_{\alpha,\beta,\gamma}\,\chi_{0}(\beta^{2}-\beta,\gamma^{2}-\gamma)=7(q+1)$.
  • $\sum_{\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}}\tau_{\alpha,\beta,\gamma}\,\big(\chi_{0}(\alpha,\beta^{2}-\beta)+\chi_{0}(\alpha-1,\beta-\gamma)\big)=6(q+1)$.
  • $\sum_{\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}}\tau_{\alpha,\beta,\gamma}\,\chi^{\times}(\alpha^{2}-\alpha)\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)=4(q-2)$.
  • $\sum_{\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}}\tau_{\alpha,\beta,\gamma}\,\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)=12.$
  • $\sum_{\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}}\tau_{\alpha,\beta,\gamma}\,\big(\chi_{0}(\alpha,\beta,\gamma-1)+\chi_{0}(\alpha,\beta-1,\gamma)\big)=6$.
  • $\sum_{\alpha,\beta,\gamma\in\mathfrak{o}/\mathfrak{m}}\tau_{\alpha,\beta,\gamma}\,\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)=4(q+1).$
We now present the calculus of $S_{i}(x_{0},x_{1},x_{2})$ and $\mathcal{W}_{i}(r_{0},r_{1},r_{2})$, and finally we express $W_{i}=W_{i}(t_{0},t_{1},t_{2})$ as a rational function of
$q,\quad X=q^{-t_{0}},\quad Y=q^{-t_{1}}\quad\text{and}\quad Z=q^{-t_{2}}.$
  • The expression for $W_{1}$ is a particular case of 9:
    $W_{1}$ $=\frac{(1+q+q^{2})(1+q)\,q^{-3}}{(1-X)(1-XY)(1-XZ)}.$
  • We have $W_{2}=\mathcal{W}_{2}(t_{0}+t_{2},t_{0}+t_{1},t_{0}+t_{1}+t_{2})$. Now
    $S(1,1,\pi x_{2},1,\alpha,\beta,\gamma)$ $=\chi_{0}\!\bigl(\beta^{2}-\beta-\alpha(\gamma^{2}-\gamma)\bigr)\,|\pi^{3}x_{2}|;$
    $S_{2}(x_{0},x_{1},x_{2})$ $=(q^{2}+8q+1)\,|\pi^{3}x_{2}|.$
    $\mathcal{W}_{2}(r_{0},r_{1},r_{2})$ $=(q^{2}+8q+1)\cdot\frac{1}{(1-q^{-1})^{3}}\int|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}\,|\pi^{3}x_{2}|\,|d(x_{0},x_{1},x_{2})|$
    $=(q^{2}+8q+1)\,q^{-3}\,\zeta_{q}(r_{0})\zeta_{q}(r_{1})\zeta_{q}(r_{2}+1).$
    Hence
    $W_{2}$ $=\frac{(q^{2}+8q+1)\,q^{-3}}{(1-XZ)(1-XY)(1-q^{-1}XYZ)}.$
  • We have $W_{3}=\mathcal{W}_{3}(t_{0}+t_{2},t_{0}+t_{1}+t_{2},t_{2})$. Now
    $S(1,1,\pi^{2}x_{1}x_{2},\pi x_{2},\alpha,\beta,\gamma)$ $=\chi_{0}(\beta^{2}-\beta,\gamma^{2}-\gamma)\,|\pi^{4}x_{1}x_{2}^{2}|;$
    $S_{3}(x_{0},x_{1},x_{2})$ $=7(q+1)\,|\pi^{4}x_{1}x_{2}^{2}|.$
    Hence
    $\mathcal{W}_{3}(r_{0},r_{1},r_{2})$ $=7(q+1)\,\frac{1}{(1-q^{-1})^{3}}\int|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}\,|\pi^{4}x_{1}x_{2}^{2}|\,|d(x_{0},x_{1},x_{2})|$
    $=7(q+1)\,q^{-4}\,\zeta_{q}(r_{0})\zeta_{q}(r_{1}+1)\zeta_{q}(r_{2}+2).$
    Hence
    $W_{3}=\frac{7(q+1)\,q^{-4}}{(1-XZ)(1-q^{-1}XYZ)(1-q^{-2}Z)}.$
For the calculation of $W_{4},\,W_{5},\,W_{7}\,$ and $W_{8}\,$, we shall need the following integral, where $t\in\mathbb{C}$ and $\epsilon\in\{0,1\}$.
$\frac{1}{1-q^{-1}}\int_{\pi^{\epsilon}a\mid x}\frac{1}{|a|}\,|x|^{t-1}\,|d(x,a)|$ $=|\pi|^{\epsilon t}\,\frac{1}{1-q^{-1}}\int|a|^{t-1}|y|^{t-1}\,|d(y,a)|=|\pi|^{\epsilon t}\,(1-q^{-1})\,\zeta_{q}(t)^{2}.$
  • We have $W_{4}=\mathcal{W}_{4}(t_{0}+t_{1},t_{0}+2t_{1}+t_{2},t_{0}+t_{1}+t_{2})$. Now
    $S(1,\pi x_{1},\pi^{2}x_{1}^{2}x_{2},1,\alpha,\beta,\gamma)$ $=\Bigl(\chi_{0}(\alpha,\beta^{2}-\beta)+\chi_{0}(\alpha-1,\beta-\gamma)+\chi^{\times}(\alpha^{2}-\alpha)\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)\Bigr)\,|\pi^{4}x_{1}^{3}x_{2}|$
    $\quad+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|\pi^{3}x_{1}^{3}x_{2}|\int_{\pi a\mid x_{1}}\frac{1}{|a|}\,|da|;$
    $S_{4}(x_{0},x_{1},x_{2})$ $=(10q-2)\,|\pi^{4}x_{1}^{3}x_{2}|+12\,|\pi^{3}x_{1}^{3}x_{2}|\int_{\pi a\mid x_{1}}\frac{1}{|a|}\,|da|.$
    $\mathcal{W}_{4}(r_{0},r_{1},r_{2})$ $=\frac{1}{(1-q^{-1})^{3}}\int|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}(10q-2)\,|\pi^{4}x_{1}^{3}x_{2}||d(x_{0},x_{1},x_{2})|$
    $\quad+\frac{1}{(1-q^{-1})^{3}}\int_{\pi a\mid x_{1}}|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}\cdot 12\,|\pi^{3}x_{1}^{3}x_{2}|\frac{1}{|a|}\,|d(x_{0},x_{1},x_{2},a)|$
    $=(10q-2)\,q^{-4}\,\zeta_{q}(r_{0})\zeta_{q}(r_{1}+3)\zeta_{q}(r_{2}+1)+12\,q^{-6-r_{1}}\,(1-q^{-1})\zeta_{q}(r_{0})\zeta_{q}(r_{1}+3)^{2}\zeta_{q}(r_{2}+1).$
    Hence,
    $W_{4}$ $=\frac{(10q-2)\,q^{-4}\,}{(1-XY)(1-q^{-3}XY^{2}Z)(1-q^{-1}XYZ)}+\frac{12\,q^{-6}\,XY^{2}Z\,(1-q^{-1})\,}{(1-XY)(1-q^{-3}XY^{2}Z)^{2}(1-q^{-1}XYZ)}.$
  • We have $W_{5}=\mathcal{W}_{5}(t_{0}+t_{1},\,t_{1},\,t_{0}+2t_{1}+t_{2})$. Note that
    $S(\pi x_{1},\pi^{2}x_{1}x_{2},\pi^{3}x_{1}x_{2}^{2},1,\alpha,\beta,\gamma)$ $=\Bigl(\chi_{0}(\alpha,\beta^{2}-\beta)+\chi_{0}(\alpha-1,\beta-\gamma)\Bigr)\,|\pi^{6}x_{1}^{2}x_{2}^{3}|$
    $\quad+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|\pi^{5}x_{1}^{2}x_{2}^{3}|\int_{a\mid x_{2}}\frac{1}{|a|}\,|da|;$
    $S_{5}(x_{0},x_{1},x_{2})$ $=6(q+1)\,|\pi^{6}x_{1}^{2}x_{2}^{3}|+12\,|\pi^{5}x_{1}^{2}x_{2}^{3}|\int_{a\mid x_{2}}\frac{1}{|a|}\,|da|.$
    $\mathcal{W}_{5}(r_{0},r_{1},r_{2})$ $=\frac{1}{(1-q^{-1})^{3}}\int|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}\,6(q+1)\,|\pi^{6}x_{1}^{2}x_{2}^{3}||d(x_{0},x_{1},x_{2})|$
    $\quad+\frac{1}{(1-q^{-1})^{3}}\int_{a\mid x_{2}}|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}\,12\,|\pi^{5}x_{1}^{2}x_{2}^{3}|\frac{1}{|a|}\,|d(a,x_{0},x_{1},x_{2})|$
    $=6(q+1)\,q^{-6}\,\zeta_{q}(r_{0})\zeta_{q}(r_{1}+2)\zeta_{q}(r_{2}+3)+12\,q^{-5}(1-q^{-1})\zeta_{q}(r_{0})\zeta_{q}(r_{1}+2)\zeta_{q}(r_{2}+3)^{2}.$
    Hence,
    $W_{5}$ $=\frac{6(q+1)\,q^{-6}\,}{(1-XY)(1-q^{-2}Y)(1-q^{-3}XY^{2}Z)}+\frac{12\,q^{-5}(1-q^{-1})\,}{(1-XY)(1-q^{-2}Y)(1-q^{-3}XY^{2}Z)^{2}}.$
  • We have $W_{6}=\mathcal{W}_{6}(t_{0}+t_{1}+t_{2},\,t_{1}+t_{2},\,t_{2})$. Note that
    $S(1,\pi x_{1},\pi^{2}x_{0}x_{1}^{2}x_{2},\pi x_{1}x_{2},\alpha,\beta,\gamma)$ $=\Bigl(\chi_{0}(\alpha,\beta,\gamma-1)+\chi_{0}(\alpha,\beta-1,\gamma)\Bigr)\,|\pi^{4}x_{0}x_{1}^{4}x_{2}^{2}|$
    $\quad+\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)\,|\pi^{4}x_{0}x_{1}^{3}x_{2}^{2}|;$
    $S_{6}(x_{0},x_{1},x_{2})$ $=6\,|\pi^{4}x_{0}x_{1}^{4}x_{2}^{2}|+4(q+1)\,|\pi^{4}x_{0}x_{1}^{3}x_{2}^{2}|.$
    $\mathcal{W}_{6}(r_{0},r_{1},r_{2})$ $=\frac{1}{(1-q^{-1})^{3}}\int|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}\left(6\,|\pi^{4}x_{0}x_{1}^{4}x_{2}^{2}|+4(q+1)\,|\pi^{4}x_{0}x_{1}^{3}x_{2}^{2}|\right)\,|d(x_{0},x_{1},x_{2})|$
    $=6\,q^{-4}\,\zeta_{q}(r_{0}+1)\zeta_{q}(r_{1}+4)\zeta_{q}(r_{2}+2)+4(q+1)\,q^{-4}\,\zeta_{q}(r_{0}+1)\zeta_{q}(r_{1}+3)\zeta_{q}(r_{2}+2).$
    Hence,
    $W_{6}$ $=\frac{6\,q^{-4}\,}{(1-q^{-1}XYZ)(1-q^{-4}YZ)(1-q^{-2}Z)}+\frac{4(q+1)\,q^{-4}\,}{(1-q^{-1}XYZ)(1-q^{-3}YZ)(1-q^{-2}Z)}.$
  • We have $W_{7}=\mathcal{W}_{7}(t_{0}+t_{1}+t_{2},t_{0}+2t_{1}+t_{2},t_{1}+t_{2})$. Now
    $S(1,\pi^{2}x_{1}x_{2},\pi^{4}x_{1}^{2}x_{2}^{2}x_{0},\pi x_{2},\alpha,\beta,\gamma)$ $=\Bigl(\chi_{0}(\alpha,\beta,\gamma-1)+\chi_{0}(\alpha,\beta-1,\gamma)\Bigr)\,|\pi^{7}x_{0}x_{1}^{3}x_{2}^{4}|$
    $\quad+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|\pi^{6}x_{0}x_{1}^{3}x_{2}^{3}|$
    $\quad+\chi^{\times}(\alpha^{2}-\alpha)\chi_{0}(\beta-\gamma,\gamma^{2}-\gamma)\,|\pi^{7}x_{0}x_{1}^{3}x_{2}^{3}|$
    $\quad+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|\pi^{6}x_{0}x_{1}^{3}x_{2}^{3}|\int_{\pi a\mid x_{1}}\frac{1}{|a|}\,|da|;$
    $S_{7}(x_{0},x_{1},x_{2})$ $=6\,|\pi^{7}x_{0}x_{1}^{3}x_{2}^{4}|+(16q-8)\,|\pi^{7}x_{0}x_{1}^{3}x_{2}^{3}|+12\,|\pi^{6}x_{0}x_{1}^{3}x_{2}^{3}|\int_{\pi a\mid x_{1}}\frac{1}{|a|}\,|da|.$
    $\mathcal{W}_{7}(r_{0},r_{1},r_{2})$ $=\frac{1}{(1-q^{-1})^{3}}\int|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}\left(6\,|\pi^{7}x_{0}x_{1}^{3}x_{2}^{4}|+(16q-8)\,|\pi^{7}x_{0}x_{1}^{3}x_{2}^{3}|\right)|d(x_{0},x_{1},x_{2})|$
    $\quad+12\,\frac{1}{(1-q^{-1})^{3}}\int_{\pi a\mid x_{1}}|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}|\pi^{6}x_{0}x_{1}^{3}x_{2}^{3}|\frac{1}{|a|}\,|d(a,x_{0},x_{1},x_{2})|$
    $=6\,q^{-7}\,\zeta_{q}(r_{0}+1)\zeta_{q}(r_{1}+3)\zeta_{q}(r_{2}+4)+(16q-8)\,q^{-7}\,\zeta_{q}(r_{0}+1)\zeta_{q}(r_{1}+3)\zeta_{q}(r_{2}+3)$
    $\quad+12\,q^{-9-r_{1}}(1-q^{-1})\zeta_{q}(r_{0}+1)\zeta_{q}(r_{1}+3)^{2}\zeta_{q}(r_{2}+3).$
    Hence,
    $W_{7}$ $=\frac{6\,q^{-7}\,}{(1-q^{-1}XYZ)(1-q^{-3}XY^{2}Z)(1-q^{-4}YZ)}+\frac{(16q-8)\,q^{-7}\,}{(1-q^{-1}XYZ)(1-q^{-3}XY^{2}Z)(1-q^{-3}YZ)}$
    $\quad+\frac{12\,q^{-9}\,XY^{2}Z\,(1-q^{-1})\,}{(1-q^{-1}XYZ)(1-q^{-3}XY^{2}Z)^{2}(1-q^{-3}YZ)}.$
  • We have $W_{8}=\mathcal{W}_{8}(t_{0}+2t_{1}+t_{2},t_{1},t_{1}+t_{2})$. Note that
    $S(\pi x_{1},\pi^{2}x_{0}x_{1}x_{2},\pi^{3}x_{0}^{2}x_{1}x_{2}^{2},\pi x_{2},\alpha,\beta,\gamma)$ $=\Bigl(\chi_{0}(\alpha,\beta,\gamma-1)+\chi_{0}(\alpha,\beta-1,\gamma)\Bigr)\,|\pi^{6}x_{0}^{3}x_{1}^{2}x_{2}^{4}|$
    $\quad+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|\pi^{5}x_{0}^{3}x_{1}^{2}x_{2}^{3}|$
    $\quad+\chi_{0}(\alpha^{2}-\alpha,\beta-\gamma,\gamma^{2}-\gamma)\,|\pi^{5}x_{0}^{3}x_{1}^{2}x_{2}^{3}|\int_{\pi a\mid x_{0}}\frac{1}{|a|}\,|da|;$
    ${S}_{8}(x_{0},x_{1},x_{2})=6\,|\pi^{6}x_{0}^{3}x_{1}^{2}x_{2}^{4}|+12\,|\pi^{5}x_{0}^{3}x_{1}^{2}x_{2}^{3}|+12\,|\pi^{5}x_{0}^{3}x_{1}^{2}x_{2}^{3}|\int_{\pi a\mid x_{0}}\frac{1}{|a|}\,|da|.$
    $\mathcal{W}_{8}(r_{0},r_{1},r_{2})$ $=\frac{1}{(1-q^{-1})^{3}}\int|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}\left(6\,|\pi^{6}x_{0}^{3}x_{1}^{2}x_{2}^{4}|+12\,|\pi^{5}x_{0}^{3}x_{1}^{2}x_{2}^{3}|\right)|d(x_{0},x_{1},x_{2})|$
    $\quad+\frac{1}{(1-q^{-1})^{3}}\int_{\pi a\mid x_{0}}12\,|x_{0}|^{r_{0}-1}|x_{1}|^{r_{1}-1}|x_{2}|^{r_{2}-1}|\pi^{5}x_{0}^{3}x_{1}^{2}x_{2}^{3}|\frac{1}{|a|}\,|d(a,x_{0},x_{1},x_{2})|$
    $=6\,q^{-6}\,\zeta_{q}(r_{0}+3)\zeta_{q}(r_{1}+2)\zeta_{q}(r_{2}+4)+12\,q^{-5}\,\zeta_{q}(r_{0}+3)\zeta_{q}(r_{1}+2)\zeta_{q}(r_{2}+3)$
    $\quad+12\,q^{-8-r_{0}}(1-q^{-1})\zeta_{q}(r_{0}+3)^{2}\zeta_{q}(r_{1}+2)\zeta_{q}(r_{2}+3).$
    Hence,
    $W_{8}$ $=\frac{6\,q^{-6}\,}{(1-q^{-3}XY^{2}Z)(1-q^{-2}Y)(1-q^{-4}YZ)}+\frac{12\,q^{-5}\,}{(1-q^{-3}XY^{2}Z)(1-q^{-2}Y)(1-q^{-3}YZ)}$
    $\quad+\frac{12\,q^{-8}\,XY^{2}Z\,(1-q^{-1})\,}{(1-q^{-3}XY^{2}Z)^{2}(1-q^{-2}Y)(1-q^{-3}YZ)}.$
Recall that our original goal was to compute $\Xi_{3,\{1,2\}}(s_{0},s_{1},s_{2})$. We defined $(t_{0},t_{1},t_{2})=(3s_{0},s_{1}-2,2s_{2}-2)$ and obtained
$q^{t_{1}+t_{2}}\,\Xi_{3,\{1,2\}}(s_{0},s_{1},s_{2})$ $=X^{2}\,W_{1}+X\,W_{2}+XZ\,W_{3}+XY\,W_{4}+XY^{2}\,W_{5}+W_{6}+XY^{2}Z\,W_{7}+Y\,W_{8}.$
Using the above expressions of $W_{1},\ldots,W_{8}$, a simple computation gives
Proposition 4.2.
Setting $X=q^{-3s_{0}},Y=q^{-s_{1}+2},Z=q^{-2s_{2}+2}$, we have
$\Xi_{3,\{1,2\}}(s_{0},s_{1},s_{2})=\frac{YZ\;P(X,Y,Z;q)}{q^{11}\,(1-X)(1-XY)(1-XZ)(1-q^{-2}Y)(1-q^{-2}Z)(1-q^{-3}YZ)(1-q^{-1}XYZ)}$
where
$P(X,Y,Z;q)=a_{0}+a_{1}X+a_{2}X^{2}+a_{3}X^{3},$
with
[…]
$\quad+(-q^{5}-2q^{4}+5q^{3}+6q^{2})Y^{3}Z^{2}+(q^{3}+2q^{2}+2q+1)Y^{3}Z^{3}.$
4.3. Application: The cotype zeta function of subalgebras of $\mathfrak{o}^{4}$
Any $\mathfrak{o}$-subalgebra of $\mathfrak{o}^{n+1}$ of finite index containing the identity $(1,\ldots,1)$ has elementary divisors $\pi^{\alpha_{1}},\ldots,\pi^{\alpha_{n}},1$ for some $\alpha_{1}\geq\cdots\geq\alpha_{n}\geq 0$. The cotype zeta function of subalgebras of $\mathfrak{o}^{n+1}$ containing the identity, as defined in [10], is the multivariable series
$\sum_{\alpha_{1}\geq\cdots\geq\alpha_{n}\geq 0}f_{n+1}(\alpha_{1},\ldots,\alpha_{n})q^{-s_{1}\alpha_{1}}\cdots q^{-s_{n}\alpha_{n}},$
where $f_{n+1}(\alpha_{1},\ldots,\alpha_{n})$ denotes the number of $\mathfrak{o}$-subalgebras of $\mathfrak{o}^{n+1}$ with elementary divisors $\pi^{\alpha_{1}},\ldots,\pi^{\alpha_{n}},1$.
This is equal to the series
$\sum_{\alpha_{1}\geq\cdots\geq\alpha_{n}\geq 0}h_{n}(\alpha_{1},\ldots,\alpha_{n})q^{-s_{1}\alpha_{1}}\cdots q^{-s_{n}\alpha_{n}},$
where $h_{n}(\alpha_{1},\ldots,\alpha_{n})$ denotes the number of $\mathfrak{o}$-submodules of $\mathfrak{o}^{n}$ that are closed under multiplication and have elementary divisors $\pi^{\alpha_{1}},\ldots,\pi^{\alpha_{n}}$.
It is easy to see that the latter series equals
$\frac{1}{1-q^{-(s_{1}+\cdots+s_{n})}}+\sum_{\emptyset\neq I\subset[n-1]}\Xi_{n,I}\!\left(\frac{s_{1}+\cdots+s_{n}}{n},\left(\frac{\sum_{i\leq\iota}s_{i}}{\iota}\right)_{\iota\in I}\right).$
In particular, the cotype zeta function of $\mathfrak{o}^{4}$ is given by
$\frac{1}{1-q^{-(s_{1}+s_{2}+s_{3})}}+\Xi_{3,\{1\}}\!\left(\frac{s_{1}+s_{2}+s_{3}}{3},s_{1}\right)+\Xi_{3,\{2\}}\!\left(\frac{s_{1}+s_{2}+s_{3}}{3},\frac{s_{1}+s_{2}}{2}\right)+\Xi_{3,\{1,2\}}\!\left(\frac{s_{1}+s_{2}+s_{3}}{3},s_{1},\frac{s_{1}+s_{2}}{2}\right).$
The formulas for $\Xi_{3,\{1\}}\!\left(\frac{s_{1}+s_{2}+s_{3}}{3},s_{1}\right)$ and $\Xi_{3,\{2\}}\!\left(\frac{s_{1}+s_{2}+s_{3}}{3},\frac{s_{1}+s_{2}}{2}\right)$ follow from Proposition 3.4:
$\Xi_{3,\{1\}}\!\left(\frac{s_{1}+s_{2}+s_{3}}{3},s_{1}\right)$ $=\frac{q^{-s_{1}}}{1-q^{-(s_{1}+s_{2}+s_{3})-s_{1}+2}}\left((1+q+q^{2})\frac{q^{-(s_{1}+s_{2}+s_{3})}}{1-q^{-(s_{1}+s_{2}+s_{3})}}+\frac{6}{1-q^{-s_{1}}}\right),$
$\Xi_{3,\{2\}}\!\left(\frac{s_{1}+s_{2}+s_{3}}{3},\frac{s_{1}+s_{2}}{2}\right)$ $=\frac{q^{-(s_{1}+s_{2})}}{1-q^{-(s_{1}+s_{2}+s_{3})-(s_{1}+s_{2})+2}}\left((1+q+q^{2})\frac{q^{-(s_{1}+s_{2}+s_{3})}}{1-q^{-(s_{1}+s_{2}+s_{3})}}+\frac{7}{1-q^{-(s_{1}+s_{2})}}\right).$
The formula for $\Xi_{3,\{1,2\}}\!\left(\frac{s_{1}+s_{2}+s_{3}}{3},s_{1},\frac{s_{1}+s_{2}}{2}\right)$ follows from Proposition 4.2.
After computing the sum of these rational functions we obtain the following result.
Proposition 4.3.
Setting $x=q^{-s_{1}},\,y=q^{-s_{2}},\,z=q^{-s_{3}}$, the cotype zeta function of subalgebras of $\mathfrak{o}^{4}$ containing the identity is
$F(x,y,z)=\frac{N(x,y,z)}{(1-x)(1-xy)(1-xyz)(1-qx^{2}y)(1-q^{2}x^{2}yz)(1-q^{2}x^{2}y^{2}z)(1-q^{3}x^{3}y^{2}z)},$
where
$N(x,y,z)=A_{0}(x,y)+A_{1}(x,y)z+A_{2}(x,y)z^{2}+A_{3}(x,y)z^{3},$
where
$A_{0}(x,y)=$ $1+5x+6xy+(3q-2)x^{2}y+(3q-4)x^{3}y-6qx^{3}y^{2}-6qx^{4}y^{2}.$
$A_{1}(x,y)=$ $(q-5)x^{2}y+(q-6)x^{2}y^{2}-(q+1)x^{3}y+(1-5q-5q^{2})x^{3}y^{2}-(q+1)x^{3}y^{3}$
[…]
$+(q+q^{2})x^{5}y^{2}+(8q+8q^{2}-7q^{3})x^{5}y^{3}+(q+q^{2})x^{5}y^{4}$
[…]
$+(q^{3}+q^{4})x^{6}y^{3}+(-7q^{2}+8q^{3}+8q^{4})x^{6}y^{4}+(q^{3}+q^{4})x^{6}y^{5}$
[…]
$-(q^{5}+q^{4})x^{8}y^{4}+(q^{5}-5q^{4}-5q^{3})x^{8}y^{5}-(q^{5}+q^{4})x^{8}y^{6}$
$+(-6q^{5}+q^{4})x^{9}y^{5}+(-5q^{5}+q^{4})x^{9}y^{6}.$
$A_{3}(x,y)=$ $-6q^{4}x^{7}y^{5}+(3q^{4}-4q^{5})x^{8}y^{6}-6q^{4}x^{8}y^{5}+(3q^{4}-2q^{5})x^{9}y^{6}$
$+6q^{5}x^{10}y^{6}+5q^{5}x^{10}y^{7}+q^{5}x^{11}y^{7}.$
Remark 4.4.
Our formula coincides with the one predicted in [3, Conjecture A.2] in the case $\mathfrak{o}=\mathbb{Z}_{p}$.